A Complete Bibliography of Publications by, and about, Richard Steven Varga

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 June 2022 Version 2.01

Title word cross-reference

(0; M) [SSV88]. 1/9 [CRV84]. $15.00 [Ano95]. 1 < λ < ∞ [ESV83s]. 2 [SSV88, SV89a, SV89b, SSV91, SSV94]. [−1, +1] [VRK91]. [−1, 1] [VRK91, VRK93]. [0, +∞] [SV76a]. [0, +∞) [MV70, MRTV71]. [0, 1] [CV93, VC92]. [0, ∞) [CMV69]. [∗ ∗ ∗w(∗ ∗ 0, +∞) ∗ ∗ ∗ w ∗ ∗ [MRTV72]. ALPS [BVW78a]. P_{n}^{g}(A) [VC82]. cos(z) [VC00, VC01, VC10]. e^{−x} [CMV69]. e^{−z} [SV75a, SSV76, SSV76a, SVN76, SSV76b]. e^{z} [ACV06, BPV93, CVW91, SV75b, SV76c, SV77c, SV78c, VC90, VC94, VCL08]. G [CV73a, CV73b, CV74]. G_{m}(w) [ESV83f, ESV83i]. H [NV84, VSM80]. J_{m}(w) [ESV83g]. k [GNV86, MSV95, NV83]. L [DV73, DV74, SV67, SV72, Tau72]. L_{p} [DV74]. Λ [COSV93, CSV94a, NRV92, CNV88]. LU [VC81, VC82]. M [NV95a, Var76a, VC81, VC82]. n [CSV80a, DKV63]. ω [EV93a, EV77, Var77]. p [LV89, LV90, Var59b, VNC84]. P_{n}^{(α_{n},β_{n})}(x) [MSV79]. £ [ESV83i, ESV83a, ESV83r]. Q_{m}(w) [ESV83f, ESV83i]. sin(z)

[VC00, VC01, VC10]. τ [Var77]. U_{m}(w) [ESV83i]. x − y [Var57a]. X^{α} [CV93, VC92]. ξ [CSV94b]. Z [NV95b]. ζ [CVV90]. |Q_{m}(w)| [ESV83c]. |U_{m}(w)| [ESV83c]. |w| = 1 [ESV83g]. |x| [VRK91, VRK93]. ||A^{−1}||_{∞} [Var76b].

-cyclic [LV89, LV90, Var59b, VNC84]. -error [DV74]. -function [CSV94b, CVV90]. -functions [ESV83i, ESV83a, ESV83r, CV73a, CV73b, CV74]. -ISBN [VR91b]. -matrices [EV77, Var77, NV95b, NV95a, VC81, NV84, VSM80, VC82]. -matrix [Var76a]. -periodic [SSV88, SV89b, SSV91, SSV94, SV89a]. -spline [DV73, DV74, SV72]. -splines [SV67]. -step [GNV86, MSV95, NV83]. -th [CSV80a].

11 [Bue66]. 13-=14 [RRV93]. 14 [GV66]. 1968 [BV70]. 1975/76 [SSV76a, SV76c]. 1976 [SV77d]. 1983 [GMSV84]. 1989 [RSSV90]. 1992 [RRV93]. 1994 [BV96a, FV94]. 1996 [Var98].

2022 [BC22]. 25 [SSV76a, SV76c]. 26 [SV77a].

3 [Bul94, VR91b]. 3-11-013784-4 [Bul94]. 3-540-12760-7 [VR91b]. 3-540-21100-4 [Kol05].

4 [Bul94].

50th [Gau94].

6 [BV59a].

7 [VR91b].

=14 [RRV93].

Abstract [Var67a]. Acceleration [ENV92]. accuracy [CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, PCV68]. Accurate [Var70a, Var70b]. Adaptive [MSV95]. ADI [VN93c]. admissible [ESV83a]. Advanced [GVC53, SGV^{+}52]. Algebra [NV94, RRV93, Var76a, Var78, Bul94]. algorithm [VN93a]. Allgower [VR91b]. Alston [Var66a, CV94]. Alternating [BVY62, DKV63, BV59b, Var60c, Var60a, Var63]. Ambikeshwar [JSV^{+}04]. American [BV70]. analiz [VV74]. analize [VV74]. analogue [Mar67]. Analysis [BV96a, Var62, Var67a, Var71a, Var00a, Var09, CR00, LSH^{+}81, NV83, POVS^{+}88, Var71b, VN93c, Var64a, Var65b, Var66a]. analytic [Ano95, GV83]. Angular [ACV06, SV76a]. Anniversary [Gau94]. Appl [GV66]. Application [HV71a, PVW66, BPV93, VN93c, VR12]. Applications [SV77d, Var67a, Var67b, BSV68, CMV69, CVV90, CRV91, JV69, PV70, Var71b, Var73b, Var78]. Applied [NdPV84, PV72a, BV70, EV93c, LV89, LV90, NV84, PV72b, Var60c, Var60a]. approach [SUV80]. approksimatsii [VV74]. approksimatsiyakh [VRK91]. Approx [VR91b]. approximants [SV75a, SV75b, SV76c, SV77c, SV78c, VC94]. approximating [SV77b]. Approximation [GMSV84, SV77d, Var71a, Var82, VCK08, APV01b, CRV84, CV93, EVN87, HV71a, KSV97, LRV75, MRTV71, MRTV72, PV97a, PV97b, PV98a, PV98b, PV99b, PV99a, RV89a, RV89b, SV78d, Var71b, Var72b, VC85, VK86, VC86, VC87, VR91a, VC92]. Approximations [PV72a, CMV69, MV70, PV70, SVN76, SV78a, Var72a, Var73a, Var73b, VRK91, VRK93]. Approximationsmethoden [EVN87]. April [BV70]. arguments [Var76b]. Arnoldi [SV93]. Articles [CV76]. assertion [ESV83r]. Associated [Var73c, CRV91, ESV83i, EV93d]. Asymptotically [KV72]. Asymptotics [CVW91, VC90, VC94]. August [Gau94]. Austin [KH90].

behavior [SV76b]. Berlin [Bul94, Kol05]. Bernstein [Var68b, VC85, VK86, VC86, VC87]. Bernsteina [VC86]. Besov [HV71a]. Bessel [dBSV80, dBSV81a, dBSV81b]. Best [EV93b, KV72, CV93, EV93a, SV77b, SV78a, VRK91, VR91a, VC92, VRK93]. between [Var73c]. Biehler [CV89a]. biharmonic [RV88]. Biostereometric [POVS^{+}88, LSH^{+}81, SPL^{+}86]. Birkhoff [Var98, CSV80a]. Birthday [Cia98, FV94, AEF^{+}10]. Block [CV59, FV62, HV64, NdPV84]. Book [Ano95, Bul94, Fix93, Kol05, Var66a, Var67b, VR91b]. both [LSV79a]. bound [CNV88, COSV93, NRV92]. boundaries [VK95]. Boundary [PV96, CPV69, CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, CV71, HSV69, HV71b, JV69, PPV69, Var66b, Var70a, Var70b]. bounded [ORV88]. bounding [Var76b]. Bounds [LSV79a, DV73, DV74, GV83, MV69, NV80, PV72b, PV70, RTV89, SV72, Var69]. Brauer [Var01a]. Breast [LSH^{+}81, POVS^{+}88, SPL^{+}86]. British [Gau94]. Brualdi [BBCH08, Var01a]. Bruijn [CNV88, COSV93, CSV94a, NRV92]. Bueckner [GV66].

calculations [Var92a]. can [Var92a]. cancer [LSH^{+}81, POVS^{+}88]. capillaries [GPMV91]. cardinality [VC82]. Carolina [BV70]. Cartesian [KVC12]. case [BNV82, EV93c, SSV88]. Cassini [VK99, Var01a]. Cauchy [BV65]. century [Gau94]. certain [MV70, MRTV71, MRTV72, SV76e, SV77a]. changes [GVB58]. chaotic [VR12]. Chapters [Var92c, Var92d]. Chebyshev [CCV72, CMV69, GTV90, GV61a, GV61b, MV70, MRTV71, MRTV72, RV78, RV89a, RV89b, Var57b, Var72a, Var73a]. Chislennie [VRK91]. chislennom [VV74]. circle [FV62]. Circles [Var04, Kol05]. circulant [Var54a]. circular [ESV83b]. circumference [ESV83g]. Class [NV95a, CV89b]. Classes [NV95b]. Co [Bul94]. Collection [CV76]. collections [CV74]. Columbia [Gau94]. Commentaries [CGO07]. Comments [GV66]. Comparison [Var57b, CCV72]. Comparisons [CV84]. completely [Var68a]. complex [APV01b, BPV93, BV81, EVN87, NV12, PV97b, PV98a, PV98b, PV99b, PV99a, RV89a, RV89b, SV77b, SV78a, SV82, SV83b, VR91a, VCK08]. Computation [CGO07, RRV93, Var86, Var90b, CR00, Gau94, Var88, Var92c, Var92d]. Computational [VR91b, Gau94, Kol05, RSSV90]. Computations [Fix93, CRV84, Var00c]. Computing [Nas90]. concentration [VP96]. concerning [SV76e, SV77a, SV77e]. conditions [CSV68b, CSV68d, CV90]. conduction [CMV69]. Conference [GMSV84, RSSV90, BV96a, RRV93]. Congratulations [FV94]. conjecture [CRV84, KV95, VC85, VK86, VC86, VC87]. Conjectures [Var86, Var90b, SV80, Var88, Var92c, Var92d, dBSV80, Fix93]. Connection [Var73c]. connections [VSM80]. constant [CNV88, COSV93, CSV94a, NRV92]. continua [APV01b]. Continuation [VR91b]. continuum [BV70]. Contour [GTV90]. convection [PVW66, PCV68]. conventions [ESV83k]. Convergence [NdPV84, PV72a, RV88, SV75a, AV76, HN88, NV84, PV72b, SSV76, SSV76a, SVN76, SSV76b, SV83b, SSV94]. converging [SV76a]. Converse [CSV85a, MRTV71, MRTV72]. convex [APV01a]. counterexample [VP96]. Crank [CCV72]. Criticality [BV58, BV59a]. Crude [ESV83c]. curve [ESV83b, ESV83h, PV97a, PV98b]. cyclic [LV89, LV90, Var59b, VNC84]. cyclically [HV64]. cylindrical [GPMV91].

D [GV66]. D. [Bue66]. December [GMSV84, SV77d]. Decision [Ano95]. decomposition [FKSV07, SV06]. Dedicated [CV76, AEF^{+}10]. defects [GPMV91]. degrees [VP96]. Dependence [HV70]. dependent [CV71, VM59]. detection [LSH^{+}81, POVS^{+}88]. Determinants [Tau49]. determining [VN93a]. diagonal [BVW78a, BVW78b, Var76b, Var76c]. Diagonally [NV94, FV62, VN93a]. differences [SSV81]. differential [BSV68, ESV83d, Var61b, Var72b]. Diffusion [Var67b, PVW66, PCV68, Var57a, VM59, Var61a]. diffusion-convection [PVW66]. dimensional [CSV67, CV70, HV71b, Mar67, Var61a]. Direction [DKV63, BV59b, BVY62, Var60c, Var60a, Var63]. discrepancy [VCL08]. Discrete [CV70, Var66c, Var72a, Var73a]. Discretization [BV65]. Discussion [ESV83e]. Disks [MV68a, MV68b, Var64b, Var01a]. distribution [ACV06, PV97a, PV98b]. divergence [NV84]. DM [Bul94, VR91b]. domain [APV01a, HN88, SV83b]. domains [EV93d, NV84, SV76d]. dominance [Var76b, Var76c]. Dominant [NV94, FV62, VN93a]. Dr. [KVW10]. due [HN88]. Durham [BV70]. during [GVB58]. dynamical [VCL08].

economic [Ano95]. Editor [FV94]. Editorial [Cia98]. Eds [Bul94, VR91b]. effect [HSV69, HV71b]. Eigenfunctions [PV72a, PV72b]. Eigenvalue [PV72a, CKV04, JV69, PV72b, Var01c, CSV68c]. Eigenvalues [HVW70, MV68a, MV68b, Var54a, Var73c, KCV09, KVC12, Var64b]. eightieth [AEF^{+}10]. Elasticity [Bue66, GV63]. electrostatics [SUV80]. element [Var73b]. Elliptic [Var67b]. endpoints [LSV79a]. energy [GPMV91]. Eneström [ASV79, ASV81, VW85]. Entire [Var54b, CSV92, KSV97, MV70, MRTV71, MRTV72]. entries [BVW78a, BVW78b]. environments [Ano95]. equation [ESV83d, RV88, Var57a]. Equations [SV93, Var65a, Var67b, BSV68, CCV72, ENV85a, ENV85b, EVN87, GNV86, HV64, MSV95, PVW66, PCV68, VM59, Var61b, Var61a, Var72b]. equiconvergence [BSV94a, BSV94b]. equidistant [CSV80c, SV89b, SSV91, SSV94]. Errata [BV59a, SSV76a]. Erratum [SV76c, SV77a]. Error [GV83, PV70, SV72, Var69, DV73, DV74, PV72b]. errors [BV65, HSV69, HV71b]. estimated [FKSV07]. Estimates [ESV83f, ESV83g, PV72a, ESV83c, RV88]. Eugene [Var67b]. Exact [NV84]. Existence [ESV83h]. exponential [SV76b]. Extended [CRV84, DV73, DV74]. extension [ASV81, BSV94a, BSV94b, CSV80b, SSV81, Var68b]. Extensions [Var73b, MRTV71, MRTV72, SV83a]. extreme [EV93d].

Faber [BSV94b, EV93d, PV96, SV93]. Factorization [Var60b, VC81, VC82]. factorizations [VSM80]. Families [VL73]. Federico [RSSV90]. Fejér [ORV88]. female [SPL^{+}86]. field [BV70]. figures [VR91b]. finite [Var73b]. Florida [GMSV84, SV77d]. Fourier [CV89a, GV56, PV96]. France [BV96a]. Frankel [Var59b]. Free [SV76f, GPMV91, SV76e, SV77a, Var52]. Frobenius [NV12]. Function [RSSV90, CV70, CVV90, CSV94b, ESV83s, VRK91, VRK93, VK95]. Functional [Var71a, VW85]. Functions [Var54b, APV01b, CV73a, CV73b, CV74, CSV92, ESV83i, ESV83a, ESV83r, GV83, JV69, KSV97, MV70, MRTV71, MRTV72, SSV76, SV76a, SSV76a, SSV76b, SV76d, SV77b, SV77e, SV78a, SSV81, Var68a]. funktsii [VRK91]. Funktsional′ [VV74].

G [Ano95, SV80]. G. [SV80]. Galerkin [CPV69, PV72a, PV70]. Garrett [Var98]. Gatlinburg [Var90a]. Gauss [GTV90]. Gaussian [GV83]. Gene [CGO07]. generalization [Var59b, Var92b]. Generalizations [HV70, JV69, FV62]. Generalized [NV95a, KCV09, KVC12, RTV89, VW85, dBSV80, dBSV81a, dBSV81b]. Gennes [GPMV91]. genus [ESV83i, ESV83a]. Geometric [SSV76, SVN76, SSV76b, SV76d, SV83b, SSV76a]. geometrically [SV76a]. geometry [Var57a]. Georg [VR91b]. German [AV76, EVN87]. Gerschgorin [FV62, HV70, HVW70, KV95, LV66a, MV68a, MV68b, Var64b, Var65c, Var70c, VL73, VK95, Var01a, Var01b]. Gersgorin [Kol05, BBCH08, EV77, FKSV07, KCV09, SV06, VK99, Var01c, Var04, VCK08]. Gersgorin-type [KCV09, VK99, Var01c]. Gershgorin [CKV04].

Gershgorin-type [CKV04]. gipoteze [VC86]. Gleichungssysteme [EVN87]. Golub [CGO07]. Graham [VR91b]. Green [CV70]. Greetings [Cia98]. Griffin [GV66, Bue66]. group [Var57a, Var61a]. Gruyter [Bul94].

H [CGO07]. half [Gau94]. half-century [Gau94]. Hardcover [VR91b]. Hardy [TV99]. having [Tau72, VP96]. Hayman [ESV83a]. heat [CMV69]. Heidelberg [Kol05]. held [BV70, GMSV84, RSSV90, SV77d]. Henrici [GV88]. Hermite [BSV68, CSV80a, CV89a, Var66b]. Hermitian [ENV92, NV89]. hierarchy [Ano95]. high [CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, Var92a]. high-order [CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70]. high-precision [Var92a]. Higher [PV72a, PV72b, PPV69, PCV68, Var61b]. higher-order [PPV69, PCV68]. History [Nas90]. Householder [Var66a, CV94]. Hybrid [ELV89, SV93]. hypocycloidal [EV93d]. hypothesis [CNV86, CV88, CV90, CRV91, CSV94a, Var00c].

if [VN93a]. II [BV81, CV73b, CSV68b, CV70, GV61b, KV95, LV66a, MV68a, NRV92, PV72b, SV77c, SV81a, SSV94, SV83b, VC90, VC01, dBSV81a, dBSV81b]. III [BBCH08, BNV82, CSV68c, MV68b, SV78c, VK95, VC10]. im [EVN87]. Implicit [BV59b, BVY62, Var60c, Var60a, Var61b, Var63]. Improved [PV72b]. Inclusion [HVW70, CKV04, Var01c]. Incomplete [SUV80, VSM80, LSV79a, SV78b, SV78d, SV79, SV81a, SV81b]. increasing [LRV75]. independent [VM59]. Industrial [BV70]. Inequalities [LSV79b, CNV86, CV88, CRV91, CSV92]. inequality [RTV89, TV99, Var92b]. Infima [Var73c]. infinite [Mar67, SVN76, Var52]. Integral [CV89b, GTV90]. International [RSSV90, BV96a, SV77d]. interpolating [SSV81, SV82, SV83b]. Interpolation [CSV80b, GMSV84, SV73, Var82, APV01b, BSV68, CSV80a, CSV80c, DV73, DV74, SV83a, SSV88, SV89a, SV89b, SSV91, SSV94, SV72, Var66b, Var69, Var72b]. interpolation-type [Var66b]. Interpolatory [KRV74]. interval [LSV79a, LRV75, RV89a, RV89b, VRK91, VRK93]. intervals [SV78a]. Introduction [ESV83j, VR91b]. Inverse [NV94, NV95b, NV95a, ACV06, VN93a]. inversion [GV56]. ISBN [Bul94, Kol05, VR91b]. ISBN-10 [Kol05]. Isolated [MV68a, MV68b, Var64b]. Iteration [CV59, DKV63, EV93b, EV93a, NV80, EVN87]. Iterationsverfahren [EVN87]. Iterative [ELV89, GV61a, GV61b, NdPV84, SV93, Var57b, Var62, Var65a, Var00a, Var09, ENV85a, EV93c, GNV86, HV64, LV89, LV90, MSV95, NV84, NV83, RV88, Var60b, Var84, VN93c, KH90, Var67b]. IV [CSV68d].

J [BV59a, Bue66, GV66, MV09, VR91b]. J. [CSV80b, CSV85b, SSV81, SV83a]. Jacobi [MSV79]. January [Var98]. Jensen [CVV90, RTV89]. Josef [FV94]. journey [CR00]. Jr. [KVW10].

June [FV94].

Kakeya [ASV79, ASV81, VW85]. Kent [RRV93]. Kingdom [GMSV84]. Komplexen [EVN87]. Konvergenz [AV76, AV76]. Kopal [Var64a].

L [Ano95, BVW78b, Bul94, CSV80b, CSV85b, SSV81, SV83a, Var67b, VR91b]. Lacunary [CSV80c, SV73, SV81b, SSV88, SV89a, SV89b]. Laguerre [CV89b, CRV91, CSV92]. Landaude [GPMV91]. large [KH90, Var84]. Least [NdPV84]. Least-squares [NdPV84]. Leffler [ESV83s]. Lehmer [COSV93, CSV94b, CSV94a]. lemma [Mar67]. Level [CSV92]. life [KVW10]. line [KSV97]. Linear [ENV92, NV94, SV93, Bul94, CCV72, CSV68a, ENV85a, ENV85b, KH90, MSV95, NV83, NV89, RRV93, Var72a, Var73a, Var76a, Var78, Var84]. local [EV93d]. Localization [KVC12]. localizations [KCV09]. logarithmic [ACV06]. Lorentz [SV79, SV80]. low [VP96]. lower [CNV88, COSV93, NRV92, RTV89]. Luis [Ano95]. Luminy [BV96a].

M [KVW10]. Macon [Var65b]. Mahler [Var92b]. making [Ano95]. March [GV88, RRV93, RSSV90]. Marí [RSSV90]. Math [GV66, SV77a]. Math. [SSV76a, SV76c]. Mathematical [BV70, CR00, Fix93, Var86, Var90b, Var88, Var92a, Var92c, Var92d]. Mathematics [VR91b, BV70, Gau94, Kol05, Gau94]. Mathematik [Var90a]. Matrices [BVW78b, BV58, BV59a, HVW70, NV95b, NV95a, Var66a, Var70c, VC81, BVW78a, CV84, EV77, FV62, HV58, KV95, LV89, LV90, NV84, NV80, NV12, PVW66, Tau72, TV99, Var54a, Var59b, Var77, VSM80, VC82, VNC84, VN93b, VK95, VR12]. Matrix [CGO07, NV94, Var62, Var65a, Var00a, Var09, CR00, HV64, Var76a, VN93a, VCK08]. Maximum [Var66c]. measurement [SPL^{+}86]. mechanics [PV70]. memoriam [CV94, JSV^{+}04, MV09, Var98]. Method [CV59, EV93b, PV72a, SV93, Var57b, ESV83l, EV93a, NV84, PV72b, VNC84, VN93c, AV76]. Methods [DKV63, ELV89, ENV92, GV61a, GV61b, NdPV84, RSSV90, Var57b, Var65a, VR91b, BV59b, BVY62, CPV69, CCV72, CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, CV71, ENV85a, ENV85b, EVN87, EV93c, GNV86, HV64, KH90, LV89, LV90, MSV95, NV83, NV89, PPV69, PCV68, Var60b, Var60c, Var60a, Var61b, Var61a, Var63, Var66b, Var70a, Var70b, Var72b, Var84, VN93c]. Michigan [Var90a]. Milestones [CGO07]. Minimal [CV73a, CV73b, EV77, KV95, LV66a, Var65c, Var70c, VL73, VK95, VCK08]. minimization [GPMV91]. minimum [RV84, RV86]. Mittag [ESV83s]. Mittag-Leffler [ESV83s]. mixed [EV93c]. moduli [RV84, RV86]. Moebius [GV56]. Moment [CV88]. Monotone [CSV69]. monotonic [Var68a]. monotonicity [VK95]. motion [VCL08]. multi [Var61a]. multi-dimensional [Var61a]. multi-group [Var61a]. multigroup [VM59].

N [Var68b]. nailuchshikh [VRK91]. Nathaniel [Var65b]. near [ESV83g]. Necessary [CV90]. negative [HV58]. Nested [MV69]. Neumaier [HN88]. Neutron [Var67b]. Newman [CNV88, COSV93, CSV94a, NRV92]. nichtsymmetrische [EVN87]. Nicolson [CCV72]. no [BV59a, Bue66, GV66, SSV76a, SV76c, SV77a]. nodes [BSV94b, CSV80c, SV89b, SSV91, SSV94]. Non [ENV92, CSV68a, HV58, NV89]. Non-Hermitian [ENV92, NV89]. non-linear [CSV68a]. non-negative [HV58]. nonlinear [CSV67, CSV68b, CSV68c, CSV68d, CSV69, CNV70, CV71, GNV86, JV69, PPV69, Var70a, Var70b, CSV68b]. Nonnegative [BV58, BV59a, VR12]. Nonnegatively [Var68a]. nonpositive [BVW78a, BVW78b]. Nonsymmetric [SV93, ENV85a, ENV85b, MSV95]. nonsymmetrical [EVN87]. Nonuniqueness [SV77b, SV78a]. Normalized [CV59, RV84, RV86, Var60b, VC94]. normally [FKSV07]. Norms [KV72, Var73c, VL73]. North [BV70]. Notational [ESV83k]. Note [GTV90, Hen77, HVW70, SV73, HN88, LV89, LV90, SV83a, Var77, VN93c]. November [Var98]. Numer [SV77a]. Numer. [SSV76a, SV76c]. Numerical [BV70, BV96a, Bue66, CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, CV71, GPMV91, GV63, PCV68, RRV93, Var57a, Var61a, Var64a, Var65b, Var66a, Var67a, Var71a, VRK91, VRK93, CRV84, CV93, CPV69, ESV83e, HSV69, HV71b, PPV69, VM59, Var70a, Var70b, Var71b, Var76a, VC92, Bul94, VR91b]. numerically [CCV72]. Numerische [Var90a]. nyĭ [VV74].

o [VRK91]. O. [LV66b]. Ob [VC86]. Obituary [GV88]. Occasion [FV94, AEF^{+}10]. October [KH90]. odnoi [VC86]. off [BVW78a, BVW78b]. off-diagonal [BVW78a, BVW78b]. Ohio [RRV93]. Olga [CV76]. omega [EV93b]. one [BSV68, VC86, CSV67, CV70]. open [SV77e, Var77]. Operator [CSV69]. Operators [KV72, Var73c]. Optimal [EV93b, EV93c, EV93a, BV81]. Order [GV61a, GV61b, PV72a, CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, ESV83s, PPV69, PV72b, PCV68, Var61b]. Orderings [Var59a]. orthogonal [APV01a, BV96a]. oscillation [PVW66]. oscillators [VR12]. otrezke [VRK91]. our [ESV83e, ESV83t]. Outline [ESV83l]. ovals [KVC12, VK99, Var01a]. Over-relaxation [Var60a]. overconvergence [CSV85a, SV76a, SV76d, SV82, SV83b, VW85]. Overrelaxation [GV61a, GV61b, Var57b, Var60c, NV84, Var59b, Var59a, Var63, Var73b, VNC84].

P [BVW78b]. PA [Ano95]. Padé [SV75a, SV75b, SV76b, SV76c, SV77d, SV77c, SV78c, VC94]. pages [Ano95, Bul94, VR91b]. pair [COSV93]. Pairs [Tau72, CSV94b, CSV94a]. Paper [Bue66, GV66]. Parabolic [SV76f, CCV72, PV70, Var61b, Var72a, Var73a]. Partial [Var54b, ACV06, BPV93, BSV68, CVW91, ORV88, Var61b, Var72b, VC90, VC00, VC01, VCL08, VC10]. particular [SV89b]. partition [VK95]. Partitioned [Var70c, KV95, VK95]. paths [BV81]. Patterns [HV70]. periodic [SSV88, SV89a, SV89b, SSV91, SSV94, CSV68d]. Perron [NV12]. perturbation [VN93c]. Peter [GV88]. petroleum [PV70]. Photo [Ano08]. Physical [Var67a]. physics [BV70, Var67b]. Piecewise [BSV68]. Pittsburgh [Ano95]. Plane [Bue66, GV63, APV01b, BV81, PV97b, PV98a, PV98b, PV99b, PV99a]. point [CSV68a, PPV69, Var66b]. points [EV93d]. poles [SSV76, SV75b, SSV76a, SV76c, SSV76b, SV77c, SV78c, VC94]. political [Ano95]. Pólya [CV89b]. Polynomial [Var82, CV93, KSV97, PV97b, PV98a, PV98b]. Polynomials [BV96a, SV76f, Var57b, APV01a, CVV90, EV93d, LSV79a, LSV79b, LRV75, MSV79, RV78, RV84, RV86, SV76e, SV77a, SV77e, SV78b, SV78d, SV79, SUV80, SSV81, SV81a, SV81b, SV82, SV83b, VW85, VP96, dBSV80, dBSV81a, dBSV81b]. portion [ESV83b]. posed [Var68a]. potential [ACV06]. Power [ESV83u]. Powers [KV72, HV58]. pp [Kol05]. precision [Var92a]. Preface [BV96b, ENV84]. prescribed [LSV79b, RV86]. Presentations [Var64a, Var65b]. priblizhenii [VC86]. Price [Bul94]. Principle [Var66c]. problem [ACV06, BPV93, CSV67, CV70, CRV91, LV66b, SV89a, SV89b]. Problems [Bue66, Fix93, GVC53, GV63, NdPV84, PV72a, SGV^{+}52, Var67a, Var90b, BV65, BV70, CPV69, CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, CMV69, CV71, HSV69, HV71b, JV69, PPV69, PV72b, PV70, SV77e, SSV91, SSV94, Var66b, Var68a, Var70a, Var70b, Var72a, Var73a, Var92c, Var92d, VK99]. Proceedings [BV96a, GMSV84, RSSV90, SV77d, BV70, RRV93]. process [Ano95, Ano95]. Professor [FV94]. projectional [Var72b]. Proof [ESV83b, ESV83p, ESV83m, ESV83n, ESV83o, ESV83q, NV94, ESV83r, ORV88]. Properties [ESV83r, ESV83s, Var54b]. property [RV78, Tau72]. Publications [Ano95, Var64a, Var65b].

Quadratic [KRV74]. Quadrature [GTV90, GV83, HSV69, HV71b]. question [Var77].

R [Bul94, GV66, VR91b]. R. [Bue66, Mar67]. radii [NV80]. radius [MV69]. rate [RV88, VW85]. Rational [PV99b, SV77d, Var82, CRV84, CCV72, CMV69, MV70, MRTV71, MRTV72, PV99a, RV89a, RV89b, SSV76, SV76a, SSV76a, SVN76, SSV76b, SV76d, SV77b, SV77e, SV78a, SSV81, VRK91, VR91a, VC92, VRK93, GMSV84]. ratsionalnikh [VRK91]. ravnomernikh [VRK91]. Rayleigh [PV72a, PV72b]. reactivity [GVB58]. Reactor [BV58, BV59a, GVB58, Var67b]. Real [RV89a, VR91a, CSV92, KSV97, RV89b, SSV76, SSV76a, SSV76b, SV78a]. Recurring [Tau49, Var76c]. reduced [HV64]. Regions [SV76f, SV76e, SV77a]. regular [CV84]. Reichel [Bul94]. related [Var00c].

Relating [PV72a]. relationship [VCL08]. Relaxation [ENV92, NV89, BV81, Var60a]. Relaxationsverfahren [AV76]. Relaxationsverfahrens [AV76]. reliability [SPL^{+}86]. Remainder [GTV90]. Remark [Bue66, GV66]. Remarks [SV80]. Reminiscences [Ano00, Var90a, Var00b]. Representation [GTV90]. research [Var92a]. reservoir [PV70]. Respective [Var54b]. response [GVB58]. result [BSV94a, BSV94b, ORV88, Var68b]. Results [PV72a, CV93, CSV85a, ESV83e, ESV83t, PV72b, Var71b, Var76a, Var78, Var84, VRK91, VC92, VRK93]. Review [Ano95, Bul94, Fix93, Kol05, Var66a, Var67b, VR91b]. revisited [CSV85b]. rezultati [VRK91]. Richard [Fix93, AEF^{+}10, Ano00, Ano08, BC22, Cia98, Kol05]. Richardson [GV61a, GV61b]. Riemann [CNV86, CV88, CVV90, CV90, CRV91, CSV94b, CSV94a, Var00c]. right [LRV75]. Ritz [PV72a, PV72b, Var66b]. Rivilin [BSV94a]. Rivlin [BSV94b, MV09]. role [Var72b]. roots [CSV80a, CSV80b, SV83a]. Rosenberg [BV79, BV81, BNV82]. Rule [GTV90]. Russian [VC86, VRK91]. Ruttan [Bul94]. RWS [Ano95].

S [AEF^{+}10, Ano00, Ano08, Bue66, Bul94, Cia98, Fix93, GV66, Kol05, Mar67, Var66a, BVW78b]. S. [Var68b, VK86, VC86, VC87]. S.O.R. [NV80]. Saaty [Ano95]. Saff [Hen77]. Santa [RSSV90]. scheme [Var59b, Var59a]. schools [Var90a]. Scientific [Nas90, RRV93, Var86, Var88, Var90b, Var92c, Var92d, Var00c, CR00, Fix93]. Scott [CV94]. Second [GV61a, GV61b]. Sections [ESV83u]. sectors [SVN76]. Selected [CGO07]. Semi [ELV89, GV61a, GV61b, Var52, Var57b, ENV85a, EV93c, Var72a, Var73a]. semi-discrete [Var72a, Var73a]. Semi-infinite [Var52]. Semi-Iterative [ELV89, GV61a, GV61b, Var57b, ENV85a, EV93c]. semidiscrete [PV70]. semiiterative [ENV85b]. sense [ESV83a]. September [BV96a, GV88]. Sequences [SV76f, SV76e, SV77a]. Series [ESV83u, VR91b, Ano95, Kol05, PV96]. Set [Var54b, BV65, CKV04, VC82, VCK08]. Sets [Var70c, VL73, CSV92, EV77, FKSV07, KV95, LV66a, SV75a, SV06, Var65c, VK95, Var01a]. Seventieth [Cia98, FV94]. several [BPV93]. Sharma [JSV^{+}04]. Sharp [RTV89]. sharpenings [ORV88]. Sharpness [ASV79, ASV81, NV80, SV76e, SV77a, SV79, SV83a, Var01c, VK95]. SIAM [BV59a, Bue66, GV66]. Simultaneous [APV01b]. Singular [FKSV07, SV06, BNV82, CNV70]. singularities [PV96]. Smallest [MV68a, MV68b, Var64b]. social [Ano95]. Society [BV70]. Solution [Bue66, GV63, SGV^{+}52, Var67b, ACV06, BV70, CPV69, HSV69, HV71b, Var57a, VM59]. Solutions [GVC53, SGV^{+}52, RV88]. Solving [Var65a, CCV72, GNV86, Var61b, Var61a, Var72b, Var84]. Some [CV93, SV77e, SSV91, SSV94, Var67a, Var71b, Var86, VC92, KSV97, NV80, SV80, Var88]. SOR [AV76, EV93a, EV93b, EV93c]. South [SV77d]. Space [DKV63]. spaces [HV71a]. Sparse [NdPV84, Var84]. Special [AEF^{+}10, Var54b]. spectral [KV95, MV69, NV80]. spline [DV73, DV74, HV71a, JV69, SV72, Var69]. Splines [KRV74, SV73, SV67]. splittings [CV84]. Springer [Kol05, VR91b]. Springer-Verlag [Kol05, VR91b]. square [VCK08]. squares [NdPV84]. SSOR [HN88, LV89, LV90, NV84, VN93c]. stable [Var61b]. Statements [ESV83t]. States [GMSV84]. Stein [BV79, BV81, BNV82]. step [GNV86, MSV95, NV83]. Steven [BC22]. Stieltjes [NV94, VN93a]. stimulate [Var92a]. Stoer [VR91b, FV94]. Strictly [NV94, VN93a]. strips [Var52]. study [ENV85a, ENV85b]. Successive [GV61a, GV61b, Var57b, NV84, Var59b, Var59a, Var63, Var73b, VNC84]. sufficient [CV90]. summability [NV83]. summer [Var90a]. Sums [Var54b, ACV06, BPV93, CVW91, ORV88, VC90, VC00, VC01, VCL08, VC10]. survey [Var84]. symmetric [AV76, NV84, VNC84, VN93b]. symmetrischen [AV76, AV76]. Symposia [Var90a]. Symposium [Gau94, SV77d, BV70]. synchronization [VR12]. Systems [ENV92, SV93, Var67b, ENV85a, ENV85b, EVN87, GNV86, KH90, MSV95, NV83, NV89, Var84]. Szego [ESV83b, ESV83h, PV97a, PV98b].

table [SV76b]. Tampa [GMSV84, SV77d]. Taussky [CV76, LV66b]. techniques [HSV69, HV71b]. technological [Ano95]. Técnica [RSSV90]. teorii [VC86]. teoriya [VV74]. Term [GTV90]. th [CSV80a]. Their [Var54b, Var78, Var01c]. theme [BBCH08]. Theodore [MV09]. Theorem [ASV79, ESV83b, ESV83r, ESV83p, ESV83m, ESV83n, ESV83o, ESV83q, HV70, HVW70, Tau49, Var01b, ASV81, CSV80b, CSV85b, CV89a, FV62, Hen77, SV79, SSV81, SV83a]. Theorems [BV79, BV81, BNV82, MRTV71, MRTV72, SV76e, SV77a, Var76c, Var01c]. Theory [CSV69, RSSV90, SV77d, Var66a, Var71a, CR00, CRV84, CSV85a, NV83, NV12, RV89b, VM59, Var71b, Var72b, Var73b, Var76a, VC85, VK86, VC86, VC87, VP96, VCL08]. Thomas [Ano95]. time [CV71, VM59]. time-dependent [CV71, VM59]. time-independent [VM59]. times [KVW10]. Todd [CV76]. Topics [Var82]. transforms [CV89a, CV89b, GV56]. transient [GVB58]. trigonometric [CSV80c, SSV88, SV89a, SV89b, SSV91, SSV94]. Turán [CNV86]. two [BSV68, CSV68a, HV71b, PPV69, RV86, Var57a, Var66b, dBSV80]. two-dimensional [HV71b]. two-group [Var57a]. two-point [CSV68a, PPV69, Var66b]. TX [KH90]. type [BV79, BV81, BNV82, CKV04, KCV09, Var66b, VK99, Var01c].

Ultrametric [NV94, NV95a, TV99, VN93a, VN93b]. unbounded [SV75a, SV76d]. unified [RV89b]. Uniform [SV78d, CV93, VRK91, VC92, VRK93]. uniqueness [ESV83h]. United [GMSV84]. unity [CSV80a, CSV80b, SV83a]. Universidad [RSSV90].

University [SV77d, Var90a]. upper [NV80]. USA [RRV93]. Using [Var57b]. USSOR [LV89, LV90]. utilizing [LSH^{+}81].

v [VV74, VC86, CSV69]. Validity [SPL^{+}86]. Valparaí [RSSV90]. value [CPV69, CSV67, CSV68b, CSV68c, CSV68d, CSV68a, CSV69, CNV70, CV71, FKSV07, HSV69, HV71b, JV69, PPV69, SV06, Var66b, Var70a, Var70b]. values [RV86]. Vancouver [Gau94]. vanishing [LSV79a]. Varga [BC22, Bul94, Fix93, Kol05, VR91b, AEF^{+}10, Ano00, Ano08, BBCH08, Bue66, Cia98, HN88, Hen77, Mar67]. Vargas [Ano95]. Variables [DKV63, BPV93, BSV68]. variants [Var63]. variational [CV70, HSV69, HV71b, Var72b]. variations [BBCH08]. varying [PV99b]. Verlag [Kol05, VR91b]. via [ACV06]. VII [Ano95, Bul94]. vindication [Var01a]. vol [Ano95, Kol05]. Volume [VR91b, AEF^{+}10]. vs [RV89a, RV89b, VR91a].

W [Bul94]. Wachspress [Var67b]. Walsh [BSV94a, BSV94b, CSV80b, CSV85a, CSV85b, SSV81, SV83a]. Weighted [KSV97, PV97b, PV98a, PV99a, PV97a, PV98b]. weights [PV99b]. well [BV65]. well-set [BV65]. Works [CGO07].

x [Kol05]. xenon [GVB58].

York [Kol05]. Young [KVW10, Var59b].

Zdenek [Var64a]. Zero [PV98b, SV76f, ESV83i, ESV83a, LSV79b, PV97a, SV76e, SV77a, SV83b]. Zero-Free [SV76f, SV76e, SV77a]. Zeros [ESV83u, EV93d, VC00, VC01, VC10, APV01a, ACV06, CVW91, COSV93, CSV94b, CSV94a, MSV79, SV75b, SV76c, SV77c, SV78c, Var52, VC90, VC94, VCL08, dBSV80, dBSV81a, dBSV81b]. Zur [AV76, AV76].

References

Andrievskii:2006:ADZ

[ACV06] Vladimir V. Andrievskii, Amos J. Carpenter, and Richard S. Varga. Angular distribution of zeros of the partial sums of e^{z} via the solution of inverse logarithmic potential problem. Computational Methods and Function Theory, 6(2):447–458, 2006. ISSN 1617-9447 (print), 2195-3724 (electronic). URL http://www.math.kent.edu/

          ~varga/pub/paper_237.pdf.

Andriyevskyy:2009:SVD

[AEF^{+}10] Volodymyr Andriyevskyy, Michael Eiermann, Roland Freund, Jing Li, Volker Mehrmann, Reinhard Nabben, Lothar Reichel, and Daniel B. Szyld. Special volume dedicated to Richard S. Varga on the occasion of his eightieth birthday. Electronic Transactions on Numerical Analysis, 36:vii, 2009/10. ISSN 1068-9613 (print), 1097-4067 (electronic).

Anonymous:1995:BRDb

[Ano95] Anonymous. Book review: Decision making in economic, political, social and technological environments with the analytic hierarchy process: the analytic hierarchy process series, vol. VII: By Thomas L. Saaty and Luis G. Vargas. RWS Publications, Pittsburgh, PA. 330 pages. $15.00. Computers and Mathematics with Applications, 29(3):116, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0898122195901094.

Anonymous:2000:RRV

[Ano00] Anonymous. Reminiscences of Richard S. Varga. Numerical Algorithms, 25(1–4):15–16, September 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/5058/
3/1/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/3/1/fulltext.pdf.

Anonymous:2008:PRV

[Ano08] Anonymous. Photo of Richard S. Varga. Linear Algebra and its Applications, 429(10):x, November 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Andrievskii:2001:ZPO

[APV01a] V. V. Andrievskii, I. E. Pritsker, and R. S. Varga. On zeros of polynomials orthogonal over a convex domain. Constructive Approximation, 17(2):209–225, 2001. ISSN 0176-4276 (print), 1432-

0940 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_228.pdf.

Andrievskii:2001:SAI

[APV01b] Vladimir V. Andrievskii, Igor E. Pritsker, and Richard S. Varga. Simultaneous approximation and interpolation of functions on continua in the complex plane. Journal de Mathématiques Pures et Appliquées, 80(4):373–388, 2001. CODEN JMPAAM. ISSN 0021-7824 (print), 1776-3371 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_231.pdf.

Anderson:1979:EKT

[ASV79] N. Anderson, E. B. Saff, and R. S. Varga. On the Eneström– Kakeya theorem and its sharpness. Linear Algebra and its Applications, 28(??):5–16, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_116.pdf;   http://
www.sciencedirect.com/science/article/pii/0024379579901137.

Anderson:1981:EEK

[ASV81] N. Anderson, E. B. Saff, and R. S. Varga. An extension of the Eneström–Kakeya theorem and its sharpness. SIAM Journal on Mathematical Analysis, 12(1):10–22, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_125.pdf.

Alefeld:1976:KSR

[AV76] G. Alefeld and R. S. Varga. Zur Konvergenz des symmetrischen Relaxationsverfahren. (German) [On the convergence of the symmetric SOR method]. Numerische Mathematik, 25(3):291–295, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_93.pdf.

Alefeld:1975:KSR

[AV76] G. Alefeld and R. S. Varga. Zur Konvergenz des symmetrischen Relaxationsverfahrens. Numerische Mathematik, 25(3):291–295, 1975/76. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Boros:2008:GVI

[BBCH08] Endre Boros, Richard A. Brualdi, Yves Crama, and A. J. Hoffman. Geršgorin variations III: On a theme of Brualdi and Varga. Linear Algebra and its Applications, 428(1):14–19, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brezinski:2022:RSV

[BC22] Claude Brezinski and Philippe G. Ciarlet. Richard Steven Varga, 1928–2022. Numerical Algorithms, 90(1):1–2, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s11075-022-01292-0.

Buoni:1982:TSR

[BNV82] John J. Buoni, Michael Neumann, and Richard S. Varga. Theorems of Stein–Rosenberg type. III. The singular case. Linear Algebra and its Applications, 42(??):183–198, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_134.pdf;   http://www.sciencedirect.com/science/
article/pii/0024379582901483.

Barnard:1993:APS

[BPV93] Roger W. Barnard, Kent Pearce, and Richard S. Varga. An application from partial sums of e^{z} to a problem in several complex variables. Journal of Computational and Applied Mathematics, 46(1–2):271–279, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_201.pdf;   http://
www.sciencedirect.com/science/article/pii/037704279390301Q.

Computational complex analysis.

Birkhoff:1968:PHI

[BSV68] G. Birkhoff, M. H. Schultz, and R. S. Varga. Piecewise Hermite interpolation in one and two variables with applications to partial differential equations. Numerische Mathematik, 11(3):232–256, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_37.pdf.

Bruck:1994:ERRa

[BSV94a] R. Brück, A. Sharma, and R. S. Varga. An extension of a result of Rivilin on Walsh equiconvergence. In H. P. Dikshit and C. A. Micchelli, editors, Advances in computational mathematics (New Delhi, India, 1993), volume 4 of Ser. Approx. Decompos., pages 225–234. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. URL http://www.math.kent.

edu/~varga/pub/paper_208.pdf.

Bruck:1994:ERRb

[BSV94b] R. Brück, A. Sharma, and R. S. Varga. An extension of a result of Rivlin on Walsh equiconvergence (Faber nodes). In R. Zahar, editor, Approximation and computation (West Lafayette, IN, 1993), volume 119 of Internat. Ser. Numer. Math., pages 41–66. Birkhäuser Boston Inc., Cambridge, MA, USA, 1994. URL http:

//www.math.kent.edu/~varga/pub/paper_210.pdf.

Bueckner:1966:RPD

[Bue66] Hans F. Bueckner. Remark on a paper by D. S. Griffin and R. S. Varga “Numerical Solution of Plane Elasticity Problems” [J. SIAM 11 (1963), no. 4, 1046–1062]. SIAM Journal on Applied Mathematics, 14(2):414, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [GV63].

Bultheel:1994:BRN

[Bul94] A. Bultheel. Book review: Numerical linear algebra: L. Reichel, A. Ruttan, R. S. Varga (Eds.): W. de Gruyter and Co., Berlin, 1993, vii + 199 pages, Price: DM 168, ISBN 3-11-013784-4. Journal of Computational and Applied Mathematics, 50(1–3):N2– N3, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0377042794903379.

Birkhoff:1958:RCN

[BV58] Garrett Birkhoff and Richard S. Varga. Reactor criticality and nonnegative matrices. Journal of the Society for Industrial and Applied Mathematics, 6(4):354–377, December 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_8.pdf. See errata [BV59a].

Birkhoff:1959:ERC

[BV59a] Garrett Birkhoff and Richard S. Varga. Errata: “Reactor Criticality and Nonnegative Matrices” [J. SIAM 6 (1958), no. 4, 354–377]. Journal of the Society for Industrial and Applied Mathematics, 7(3): 343, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic). See [BV58].

Birkhoff:1959:IAD

[BV59b] Garrett Birkhoff and Richard S. Varga. Implicit alternating direction methods. Transactions of the American Mathematical Society, 92(1):13–24, July/September 1959. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_10.pdf.

Birkhoff:1965:DEW

[BV65] Garrett Birkhoff and Richard S. Varga. Discretization errors for well-set Cauchy problems. I. Journal of Mathematical Physics, 44: 1–23, 1965. CODEN JMAPAQ. ISSN 0097-1421. URL http://

www.math.kent.edu/~varga/pub/paper_28.pdf.

Birkhoff:1970:NSF

[BV70] Garrett Birkhoff and Richard S. Varga, editors. Numerical solution of field problems in continuum physics: proceedings of a Symposium in applied mathematics of the American Mathematical Society and the Society for Industrial and Applied Mathematics held in Durham, North Carolina, April 5–6, 1968, volume 2 of SIAM-AMS proceedings. American Mathematical Society, Providence, RI, USA, 1970. ISBN 0-8218-1321-0. LCCN QA1 .A5217 n.s. v.2 1968; QC174.45.S96 1968.

Buoni:1979:TSR

[BV79] John J. Buoni and Richard S. Varga. Theorems of Stein–Rosenberg type. In ????, editor, Numerical mathematics = Numerische Mathematik: Symposium on the occasion of the retirement of Lothar Collatz at the Institute for Applied Mathematics, University of Hamburg, January 25–26, 1979, volume 49 of Internat. Ser. Numer. Math., pages 65–75. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1979. URL http://www.math.kent.

edu/~varga/pub/paper_115.pdf.

Buoni:1981:TSR

[BV81] John J. Buoni and Richard S. Varga. Theorems of Stein–Rosenberg type. II. Optimal paths of relaxation in the complex plane. In Martin H. Schultz, editor, Elliptic Problem Solvers, pages 231–240. Academic Press, New York, NY, USA, 1981. URL http://www.

math.kent.edu/~varga/pub/paper_131.pdf.

Brezinski:1996:OPN

[BV96a] C. Brezinski and R. S. Varga, editors. Orthogonal Polynomials and Numerical Analysis. Proceedings of the international conference, Luminy, France, September 5–10, 1994, volume 11(1–4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, March 1996. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ????

Brezinski:1996:P

[BV96b] Claude Brezinski and Richard S. Varga. Preface. Numerical Algorithms, 11(1-4):IX, 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Berman:1978:MND

[BVW78a] A. Berman, R. S. Varga, and R. C. Ward. ALPS: matrices with nonpositive off-diagonal entries. Linear Algebra and its Applications, 21(3):233–244, 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:1978:PMN

[BVW78b] A. Berman, R. S. Varga, and R. C. Ward. A L P S: Matrices with nonpositive off-diagonal entries. Linear Algebra and its Applications, 21(3):233–244, September 1978. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_108.pdf;   http://
www.sciencedirect.com/science/article/pii/002437957890085X.

Birkhoff:1962:ADI

[BVY62] Garrett Birkhoff, Richard S. Varga, and David Young. Alternating direction implicit methods. In Franz L. Alt, editor, Advances in Computers, volume 3, pages 189–273. Academic Press, New York, NY, USA, 1962. ISSN 0065-2458. URL http://www.math.kent.

edu/~varga/pub/paper_20.pdf.

Cavendish:1972:CCN

[CCV72] J. C. Cavendish, W. E. Culham, and R. S. Varga. A comparison of Crank–Nicolson and Chebyshev rational methods for numerically solving linear parabolic equations. Journal of Computational Physics, 10(2):354–368, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
72.pdf;   http://www.sciencedirect.com/science/article/
pii/0021999172900721.

Chan:2007:MMC

[CGO07] Raymond H. Chan, Chen Greif, and Dianne P. O’Leary, editors. Milestones in Matrix Computation: the Selected Works of Gene H. Golub with Commentaries. Oxford University Press, Oxford, UK, 2007. ISBN 0-19-920681-3. xi + 565 + 3 pp. LCCN QA188 .G67

2007. URL http://www.loc.gov/catdir/enhancements/fy0737/
2007276086-d.html.

Ciarlet:1998:ERV

[Cia98] Philippe Ciarlet. Editorial: [to Richard S. Varga, seventieth birthday greetings]. Numerische Mathematik, 80(4):495, October 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See their earliest joint paper [CSV67].

Cvetkovic:2004:NGT

[CKV04] Ljiljana Cvetkovic, Vladimir Kostic, and Richard S. Varga. A new Geršhgorin-type eigenvalue inclusion set. Electronic Transactions on Numerical Analysis, 18:73–80, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.mcs.

kent.edu/vol.18.2004/pp73-80.dir/pp73-80.pdf.

Cody:1969:CRA

[CMV69] W. J. Cody, G. Meinardus, and R. S. Varga. Chebyshev rational approximations to e^{−x} on [0, ∞) and applications to heat conduction problems. Journal of Approximation Theory, 2(1):50–65, March 1969. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-

0430 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_50.pdf.

Ciarlet:1970:NMH

[CNV70] P. G. Ciarlet, F. Natterer, and R. S. Varga. Numerical methods of high-order accuracy for singular nonlinear boundary value problems. Numerische Mathematik, 15(2):87–99, June 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_57.pdf.

Csordas:1986:RHT

[CNV86] George Csordas, Timothy S. Norfolk, and Richard S. Varga. The Riemann hypothesis and the Turán inequalities. Transactions of the American Mathematical Society, 296(2):521–541, 1986. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_157.pdf.

Csordas:1988:LBB

[CNV88] G. Csordas, T. S. Norfolk, and R. S. Varga. A lower bound for the de Bruijn–Newman constant Λ. Numerische Mathematik, 52(5):483– 497, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_162.pdf.

Csordas:1993:NLP

[COSV93] G. Csordas, A. M. Odlyzko, W. Smith, and R. S. Varga. A new Lehmer pair of zeros and a new lower bound for the de Bruijn– Newman constant Λ. Electronic Transactions on Numerical Analysis, 1:104–111, December 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.mcs.kent.edu/

vol.1.1993/pp104-111.dir/pp104-111.pdf.

Cavendish:1969:GMN

[CPV69] J. C. Cavendish, H. S. Price, and R. S. Varga. Galerkin methods for the numerical solution of boundary value problems. Society of Petroleum Engineers journal, 9(??):800–815, June 1969. CODEN

SPTJAJ. ISSN 0197-7520. URL http://www.math.kent.edu/~
varga/pub/paper_51.pdf.

Calvetti:2000:MJT

[CR00] Daniela Calvetti and Lothar Reichel, editors. Mathematical journey through analysis, matrix theory and scientific computation, volume 25(1–4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 2000. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ???? Papers from the conference held in honor of Richard S. Varga on the occasion of his 70th birthday at Kent State University, Kent, OH, March 25–27, 1999.

Carpenter:1984:ENC

[CRV84] A. J. Carpenter, A. Ruttan, and R. S. Varga. Extended numerical computations on the “1/9” conjecture in rational approximation theory. In P. R. Graves-Morris, E. B. Saff, and R. S. Varga, editors, Rational approximation and interpolation (Tampa, Fla., 1983), volume 1105 of Lecture Notes in Math., pages 383–411. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

1984. URL http://www.math.kent.edu/~varga/pub/paper_148.
pdf.

Csordas:1991:LIA

[CRV91] G. Csordas, A. Ruttan, and R. S. Varga. The Laguerre inequalities with applications to a problem associated with the Riemann hypothesis. Numerical Algorithms, 1(3):305–329, September 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.math.kent.edu/~varga/pub/paper_185.pdf.

Ciarlet:1967:NMH

[CSV67] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. I. One dimensional problem. Numerische Mathematik, 9(5):394–430, April 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/

paper_35.pdf. See birthday greetings [Cia98].

Ciarlet:1968:NMHe

[CSV68a] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for non-linear two-point boundary value problems. In ????, editor, Programmation en Mathématiques Numériques (Actes Colloq. Internat. C.N.R.S. No. 165, Besançon, 7–14 septembre 1966), pages 217–225. Éditions du Centre National de la Recherche Scientifique (CNRS), Paris, France, 1968. LCCN

QA264 .C63 1966. URL http://www.math.kent.edu/~varga/
pub/paper_42.pdf.

Ciarlet:1968:NMHb

[CSV68b] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. II. Nonlinear boundary conditions. Numerische Mathematik, 11 (4):331–345, May 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_39.pdf.

Ciarlet:1968:NMHc

[CSV68c] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. III. Eigenvalue problems. Numerische Mathematik, 12(2):120–133, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_41.pdf.

Ciarlet:1968:NMHd

[CSV68d] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. IV. Periodic boundary conditions. Numerische Mathematik, 12(4): 266–279, November 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_45.pdf.

Ciarlet:1969:NMH

[CSV69] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. V: Monotone operator theory. Numerische Mathematik, 13(1):51–77, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_48.pdf.

Cavaretta:1980:HBI

[CSV80a] A. S. Cavaretta, Jr., A. Sharma, and R. S. Varga. Hermite–Birkhoff interpolation in the n-th roots of unity. Transactions of the American Mathematical Society, 259(2):621–628, 1980. CODEN TAM-TAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_118.pdf.

Cavaretta:1980:IRU

[CSV80b] A. S. Cavaretta, Jr., A. Sharma, and R. S. Varga. Interpolation in the roots of unity: an extension of a theorem of J. L. Walsh. Results in Mathematics = Resultate der Mathematik, 3(2):155–191, 1980.

ISSN 0378-6218. URL http://www.math.kent.edu/~varga/pub/
paper_124.pdf.

Cavaretta:1980:LTI

[CSV80c] A. S. Cavaretta, Jr., A. Sharma, and R. S. Varga. Lacunary trigonometric interpolation on equidistant nodes. In R. A. DeVore and K. Scherer, editors, Quantitative approximation (Proc. Internat. Sympos., Bonn, 1979), pages 63–80. Academic Press, New York,

NY, USA, 1980. URL http://www.math.kent.edu/~varga/pub/
paper_120.pdf.

Cavaretta:1985:CRW

[CSV85a] A. S. Cavaretta, Jr., A. Sharma, and R. S. Varga. Converse results in the Walsh theory of overconvergence. Mathematical modelling and numerical analysis = Modelisation mathématique et analyse numérique: M^{2}AN, 19(4):601–609, 1985. CODEN RMMAEV. ISSN 0764-583x (print), 1290-3841 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_153.pdf.

Cavaretta:1985:TJW

[CSV85b] A. S. Cavaretta, Jr., A. Sharma, and R. S. Varga. A theorem of J. L. Walsh, revisited. Pacific Journal of Mathematics, 118(2):313–322, 1985. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http://projecteuclid.

org/euclid.pjm/1102706441;   http://www.math.kent.edu/~
varga/pub/paper_149.pdf.

Csordas:1992:LSR

[CSV92] George Csordas, Wayne Smith, and Richard S. Varga. Level sets of real entire functions and the Laguerre inequalities. Analysis, 12(3–4):377–402, 1992. ISSN 0174-4747. URL http:/

/www.math.kent.edu/~varga/pub/paper_192.pdf;   https://
www.degruyter.com/document/doi/10.1524/anly.1992.12.34.
377/html.

Csordas:1994:LPZb

[CSV94a] George Csordas, Wayne Smith, and Richard S. Varga. Lehmer pairs of zeros, the de Bruijn–Newman constant Λ, and the Riemann hypothesis. Constructive Approximation, 10(1):107–129, 1994. ISSN 0176-4276 (print), 1432-0940 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_206.pdf.

Csordas:1994:LPZa

[CSV94b] George Csordas, Wayne [Stewart] Smith, and Richard S. Varga. Lehmer pairs of zeros and the Riemann ξ-function. In Gautschi [Gau94], pages 553–556. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48

1994. URL http://www.ams.org/books/psapm/048/;   http://
www.math.kent.edu/~varga/pub/paper_209.pdf. See also SIAM

Review, September 1995, 37(3), p. 483.

Cuthill:1959:MNB

[CV59] Elizabeth H. Cuthill and Richard S. Varga. A method of normalized block iteration. Journal of the ACM, 6(2):236–244, April 1959. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.math.kent.edu/~varga/pub/paper_9.pdf.

Ciarlet:1970:DVG

[CV70] P. G. Ciarlet and R. S. Varga. Discrete variational Green’s function. II. One dimensional problem. Numerische Mathematik, 16(2): 115–128, November 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_59.pdf.

Culham:1971:NMT

[CV71] W. E. Culham and Richard S. Varga. Numerical methods for time-dependent, nonlinear boundary value problems. Society of Petroleum Engineers journal, 11(??):374–388, December 1971. CO-DEN SPTJAJ. ISSN 0197-7520. URL http://www.math.kent.

edu/~varga/pub/paper_66.pdf.

Carlson:1973:MF

[CV73a] David H. Carlson and Richard S. Varga. Minimal G-functions. Linear Algebra and its Applications, 6(??):97–117, ???? 1973.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
76.pdf;   http://www.sciencedirect.com/science/article/
pii/0024379573900098.

Carlson:1973:MFI

[CV73b] David H. Carlson and Richard S. Varga. Minimal G-functions. II. Linear Algebra and its Applications, 7(3):233–242, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
78.pdf;   http://www.sciencedirect.com/science/article/
pii/0024379573900426.

Carlson:1974:CF

[CV74] David H. Carlson and Richard S. Varga. On collections of G-functions. Linear Algebra and its Applications, 8(1):65–76, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print),

1873-1856 (electronic).  URL http://www.math.kent.edu/~
varga/pub/paper_83.pdf;   http://www.sciencedirect.com/
science/article/pii/0024379574900093.

Carlson:1976:CAD

[CV76] David H. Carlson and Richard S. Varga, editors. Collection of Articles Dedicated to Olga Taussky Todd. American Elsevier Publishing Co., New York, NY, USA, 1976. iv + 172

pp. URL https://ac.els-cdn.com/0024379576900501/1-s2.
0-0024379576900501-main.pdf. Linear Algebra and Appl. 13

(1976), no. 1–2.

Csordas:1984:CRS

[CV84] George Csordas and Richard S. Varga. Comparisons of regular splittings of matrices. Numerische Mathematik, 44(1):23–35, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_146.pdf.

Csordas:1988:MIR

[CV88] George Csordas and Richard S. Varga. Moment inequalities and the Riemann hypothesis. Constructive Approximation, 4(2):175– 198, 1988. ISSN 0176-4276 (print), 1432-0940 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_161.pdf.

Csordas:1989:FTH

[CV89a] G. Csordas and R. S. Varga. Fourier transforms and the Hermite– Biehler theorem. Proceedings of the American Mathematical Society, 107(3):645–652, 1989. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_175.pdf.

Csordas:1989:ITL

[CV89b] George Csordas and Richard S. Varga. Integral transforms and the Laguerre–Pólya class. Complex Variables, Theory and Application, 12(1–4):211–230, 1989. CODEN CVTADV. ISSN 0278-1077. URL

http://www.math.kent.edu/~varga/pub/paper_169.pdf.

Csordas:1990:NSC

[CV90] George Csordas and Richard S. Varga. Necessary and sufficient conditions and the Riemann hypothesis. Advances in Applied Mathematics, 11(3):328–357, 1990. ISSN 0196-8858 (print), 1090-

2074 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_180.pdf.

Carpenter:1993:SNR

[CV93] Amos J. Carpenter and Richard S. Varga. Some numerical results on best uniform polynomial approximation of X^{α} on [0, 1]. Lecture Notes in Mathematics, 1550:192–222, 1993. CODEN LN-MAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0117488/;   http://www.
math.kent.edu/~varga/pub/paper_194.pdf. Methods of approx-

imation theory in complex analysis and mathematical physics (Leningrad, 1991).

Ciarlet:1994:MAS

[CV94] Philippe G. Ciarlet and Richard S. Varga. In memoriam: Alston Scott Householder, 1904–1993. Numerische Mathematik, 68(2):187, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/4068002/40680187.htm;
 http://science.springer.de/nmee/bibs/4068002/40680187.
htm.

Csordas:1990:JPA

[CVV90] George Csordas, Richard S. Varga, and István Vincze. Jensen polynomials with applications to the Riemann ζ-function. Journal of Mathematical Analysis and Applications, 153(1):112–135, 1990. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
181.pdf.

Carpenter:1991:AZP

[CVW91] A. J. Carpenter, R. S. Varga, and J. Waldvogel. Asymptotics for the zeros of the partial sums of e^{z}. I. Rocky Mountain Journal of Mathematics, 21(1):99–120, 1991. CODEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_184.pdf. Proceedings of the U.S.-

Western Europe Regional Conference on Padé Approximants and Related Topics (Boulder, CO, 1988).

deBruin:1980:TCZ

[dBSV80] M. G. de Bruin, E. B. Saff, and R. S. Varga. On two conjectures on the zeros of generalized Bessel polynomials. In E. W. Cheney, editor, Approximation theory, III (Proc. Conf., Univ. Texas, Austin, Tex., 1980), pages 261–266. Academic Press, New York, NY, USA,

1980. URL http://www.math.kent.edu/~varga/pub/paper_121.
pdf.

deBruin:1981:ZGBa

[dBSV81a] M. G. de Bruin, E. B. Saff, and R. S. Varga. On the zeros of generalized Bessel polynomials. I, II. Koninklijke Nederlandse Akademie van Wetenschappen. Indagationes Mathematicae, 43(1): 1–13, March 20, 1981. ISSN 0019-3577, 0023-3358. URL http://

www.math.kent.edu/~varga/pub/paper_127.pdf.

deBruin:1981:ZGBb

[dBSV81b] M. G. de Bruin, E. B. Saff, and R. S. Varga. On the zeros of generalized Bessel polynomials. I, II. Koninklijke Nederlandse Akademie van Wetenschappen. Indagationes Mathematicae, 43(1): 14–25, March 20, 1981. ISSN 0019-3577, 0023-3358. URL http:/

/www.math.kent.edu/~varga/pub/paper_128.pdf.

Douglas:1963:ADI

[DKV63] J. Douglas, Jr., R. B. Kellogg, and R. S. Varga. Alternating direction iteration methods for n space variables. Mathematics of Computation, 17(83):279–282, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_23.pdf.

Demko:1973:EEB

[DV73] Stephen Demko and Richard S. Varga. Extended error bounds for spline and L-spline interpolation. In ????, editor, Approximation theory (Proc. Internat. Sympos., Univ. Texas, Austin, Tex., 1973), pages 313–318. Academic Press, New York, NY, USA, 1973. LCCN

QA221 .A6446. URL http://www.math.kent.edu/~varga/pub/
paper_82.pdf.

Demko:1974:EEB

[DV74] Stephen Demko and Richard S. Varga. Extended L_{p}-error bounds for spline and L-spline interpolation. Journal of Approximation Theory, 12(3):242–264, November 1974. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_84.pdf.

Eiermann:1989:HSI

[ELV89] M. Eiermann, X. Li, and R. S. Varga. On hybrid semi-iterative methods. SIAM Journal on Numerical Analysis, 26(1):152–168, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-

7170 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_166.pdf.

Elsner:1984:P

[ENV84] L. Elsner, W. Niethammer, and R. S. Varga. Preface. Linear Algebra and its Applications, 58(??):1, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958490199X.

Eiermann:1985:SSI

[ENV85a] M. Eiermann, W. Niethammer, and R. S. Varga. A study of semiiterative methods for nonsymmetric systems of linear equations. Numerische Mathematik, 47(4):505–533, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eiermann:1985:SSM

[ENV85b] M. Eiermann, W. Niethammer, and R. S. Varga. A study of semiiterative methods for nonsymmetric systems of linear equations. Numerische Mathematik, 47(4):505–533, December 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_152.pdf.

Eiermann:1992:ARM

[ENV92] M. Eiermann, W. Niethammer, and R. S. Varga. Acceleration of relaxation methods for non-Hermitian linear systems. SIAM Journal on Matrix Analysis and Applications, 13(3):979–991, July 1992. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_

188.pdf. Iterative methods in numerical linear algebra (Copper Mountain, CO, 1990).

Edrei:1983:FGZ

[ESV83a] Albert Edrei, Edward B. Saff, and Richard S. Varga. £-functions of genus zero are admissible in the sense of Hayman. Lecture Notes in Mathematics, 1002:91–92, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.com/

chapter/10.1007/BFb0070491.

Edrei:1983:CPS

[ESV83b] Albert Edrei, Edward B. Saff, and Richard S. Varga. The circular portion of the Szegő curve (proof of theorem 3). Lecture Notes in Mathematics, 1002:77–79, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.com/

chapter/10.1007/BFb0070487/.

Edrei:1983:CE

[ESV83c] Albert Edrei, Edward B. Saff, and Richard S. Varga. Crude estimates for |U_{m}(w)| and |Q_{m}(w)|. Lecture Notes in Mathematics, 1002:63–69, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/10.
1007/BFb0070483/.

Edrei:1983:DE

[ESV83d] Albert Edrei, Edward B. Saff, and Richard S. Varga. A differential equation. Lecture Notes in Mathematics, 1002:53–60, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070480/.

Edrei:1983:DON

[ESV83e] Albert Edrei, Edward B. Saff, and Richard S. Varga. Discussion of our numerical results. Lecture Notes in Mathematics, 1002:26–40, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/10.
1007/BFb0070475/.

Edrei:1983:E

[ESV83f] Albert Edrei, Edward B. Saff, and Richard S. Varga. Estimates for G_{m}(w) and Q_{m}(w). Lecture Notes in Mathematics, 1002:49–52, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/10.
1007/BFb0070479/.

Edrei:1983:ENC

[ESV83g] Albert Edrei, Edward B. Saff, and Richard S. Varga. Estimates for J_{m}(w) near the circumference |w| = 1. Lecture Notes in Mathematics, 1002:60–62, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/10.
1007/BFb0070481/.

Edrei:1983:EUS

[ESV83h] Albert Edrei, Edward B. Saff, and Richard S. Varga. Existence and uniqueness of the Szegő curve. Lecture Notes in Mathematics, 1002:62–63, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/10.
1007/BFb0070482.

Edrei:1983:FAF

[ESV83i] Albert Edrei, Edward B. Saff, and Richard S. Varga. The functions U_{m}(w), Q_{m}(w), G_{m}(w) associated with £-functions of genus zero. Lecture Notes in Mathematics, 1002:92–94, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070492/.

Edrei:1983:I

[ESV83j] Albert Edrei, Edward B. Saff, and Richard S. Varga. Introduction. Lecture Notes in Mathematics, 1002:1–7, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070473/.

Edrei:1983:NC

[ESV83k] Albert Edrei, Edward B. Saff, and Richard S. Varga. Notational conventions. Lecture Notes in Mathematics, 1002:43–44, 1983. CO-DEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070477.

Edrei:1983:OM

[ESV83l] Albert Edrei, Edward B. Saff, and Richard S. Varga. Outline of the method. Lecture Notes in Mathematics, 1002:41–43, 1983. CO-DEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070476/.

Edrei:1983:PTb

[ESV83m] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 1. Lecture Notes in Mathematics, 1002:70–72, 1983. CO-DEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070485/.

Edrei:1983:PTc

[ESV83n] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 2. Lecture Notes in Mathematics, 1002:72–77, 1983. CO-DEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070486/.

Edrei:1983:PTd

[ESV83o] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 4. Lecture Notes in Mathematics, 1002:80–82, 1983. CO-DEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070488/.

Edrei:1983:PTa

[ESV83p] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 5. Lecture Notes in Mathematics, 1002:70, 1983. CO-DEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070484.

Edrei:1983:PTe

[ESV83q] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 6. Lecture Notes in Mathematics, 1002:82–87, 1983. CO-DEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070489/.

Edrei:1983:PFP

[ESV83r] Albert Edrei, Edward B. Saff, and Richard S. Varga. Properties of £-functions; proof of assertion I of theorem 7. Lecture Notes in Mathematics, 1002:87–91, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.com/

chapter/10.1007/BFb0070490/.

Edrei:1983:PML

[ESV83s] Albert Edrei, Edward B. Saff, and Richard S. Varga. Properties of the Mittag-Leffler function of order 1 < λ < ∞. Lecture Notes in Mathematics, 1002:44–49, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.com/

chapter/10.1007/BFb0070478/.

Edrei:1983:SOR

[ESV83t] Albert Edrei, Edward B. Saff, and Richard S. Varga. Statements of our results. Lecture Notes in Mathematics, 1002:7–25, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070474/.

Edrei:1983:ZSP

[ESV83u] Albert Edrei, Edward B. Saff, and Richard S. Varga. Zeros of Sections of Power Series, volume 1002 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 115 pp. LCCN QA3 .L28 no. 1002.

URL http://link.springer.com/book/10.1007/BFb0070472;
 http://www.springerlink.com/content/978-3-540-40040-0.

Engel:1977:MGS

[EV77] G. M. Engel and R. S. Varga. Minimal Geršgorin sets and ωmatrices. Linear and Multilinear Algebra, 5(1):1–10, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_104.pdf.

Eiermann:1993:OBS

[EV93a] M. Eiermann and R. S. Varga. Is the optimal ω best for the SOR iteration method? Linear Algebra and its Applications, 182: 257–277, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_195.pdf.

Eiermann:1993:OOB

[EV93b] M. Eiermann and R. S. Varga. Is the optimal omega best for the SOR iteration method? Linear Algebra and its Applications, 182:257–??, March 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eiermann:1993:OSI

[EV93c] Michael Eiermann and Richard S. Varga. Optimal semi-iterative methods applied to SOR in the mixed case. In L. Reichel, A. Ruttan, and R. S. Varga, editors, Numerical linear algebra (Kent, OH, 1992), pages 47–73. Walter de Gruyter, Berlin, Germany, 1993.

URL http://www.math.kent.edu/~varga/pub/paper_198.pdf.

Eiermann:1993:ZLE

[EV93d] Michael Eiermann and Richard S. Varga. Zeros and local extreme points of Faber polynomials associated with hypocycloidal domains. Electronic Transactions on Numerical Analysis, 1:49–71, September 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (elec-

tronic). URL http://etna.mcs.kent.edu/vol.1.1993/pp49-71.
dir/pp49-71.pdf.

Eiermann:1987:ING

[EVN87] M. Eiermann, R. S. Varga, and W. Niethammer. Iterationsverfahren für nichtsymmetrische Gleichungssysteme und Approximationsmethoden im Komplexen. (German) [Iteration methods for nonsymmetrical systems of equations and approximation methods in the complex]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 89(1):1–32, 1987. CODEN JDMVA7. ISSN 0012-0456 (print), 1869-7135 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_160.pdf.

Fix:1993:BRS

[Fix93] George J. Fix. Book review: Scientific Computations on Mathematical Problems and Conjectures (Richard S. Varga). SIAM Review, 35(2):318–320, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fontes:2007:SVD

[FKSV07] Natacha Fontes, Janice Kover, Laura Smithies, and Richard S. Varga. Singular value decomposition normally estimated Geršgorin sets. Electronic Transactions on Numerical Analysis, 26: 320–329, 2007. CODEN ???? ISSN 1068-9613 (print),

1097-4067 (electronic). URL http://etna.mcs.kent.edu/vol.
26.2007/pp320-329.dir/pp320-329.pdf;   http://www.math.
kent.edu/~varga/pub/paper_238.pdf.

Feingold:1962:BDD

[FV62] David G. Feingold and Richard S. Varga. Block diagonally dominant matrices and generalizations of the Gerschgorin circle theorem. Pacific Journal of Mathematics, 12(4):1241–1250, 1962. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL

http://projecteuclid.org/euclid.pjm/1103036125;   http:
//www.math.kent.edu/~varga/pub/paper_21.pdf.

Freund:1994:CEP

[FV94] Roland W. Freund and Richard S. Varga. Congratulations [to editor Professor Josef Stoer on the occasion of his seventieth birthday, June 21, 1994]. Numerische Mathematik, 68(1):1, June 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943– 1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994. See also SIAM Review, September 1995, 37(3), p. 483.

Graves-Morris:1984:RAI

[GMSV84] Peter Russell Graves-Morris, Edward B. Saff, and Richard S. Varga, editors. Rational Approximation and Interpolation: Proceedings of the United Kingdom–United States Conference held at Tampa, Florida, December 12–16, 1983, volume 1105 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1105; QA3 .L471 no.1105; QA3.L471. DM72.00. URL http://link.springer.

com/book/10.1007/BFb0072395;   http://www.springerlink.
com/content/978-3-540-39113-5.

Gutknecht:1986:SIM

[GNV86] M. H. Gutknecht, W. Niethammer, and R. S. Varga. k-step iterative methods for solving nonlinear systems of equations. Numerische Mathematik, 48(6):699–712, June 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_156.pdf.

Gartland:1991:NML

[GPMV91] E. C. Gartland, Jr., P. Palffy-Muhoray, and R. S. Varga. Numerical minimization of the Landaude–de Gennes free energy: defects in cylindrical capillaries. Molecular Crystals and Liquid Crystals (1969–1991), 199(1):429–452, ???? 1991. CODEN MCLCA5.

ISSN 0026-8941. URL http://www.math.kent.edu/~varga/pub/
paper_183.pdf.

Gautschi:1990:NCI

[GTV90] Walter Gautschi, E. Tychopoulos, and R. S. Varga. A note on the contour integral representation of the remainder term for a Gauss– Chebyshev quadrature rule. SIAM Journal on Numerical Analysis, 27(1):219–224, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_177.pdf.

Goldberg:1956:MIF

[GV56] Richard R. Goldberg and Richard S. Varga. Moebius inversion of Fourier transforms. Duke Mathematical Journal, 23(4): 553–559, December 1956. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL http://projecteuclid.

org/euclid.dmj/1077467072;   http://www.math.kent.edu/~
varga/pub/paper_3.pdf.

Golub:1961:CSIa

[GV61a] Gene H. Golub and Richard S. Varga. Chebyshev semi-iterative methods, successive overrelaxation iterative methods, and second order Richardson iterative methods. I. Numerische Mathematik, 3 (1):147–156, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_17.pdf. Reprinted in [CGO07].

Golub:1961:CSIb

[GV61b] Gene H. Golub and Richard S. Varga. Chebyshev semi-iterative methods, successive overrelaxation iterative methods, and second order Richardson iterative methods. II. Numerische Mathematik, 3 (1):157–168, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_18.pdf.

Griffin:1963:NSP

[GV63] D. S. Griffin and R. S. Varga. Numerical solution of plane elasticity problems. Journal of the Society for Industrial and Applied Mathematics, 11(4):1046–1062, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_24.pdf. See remark [Bue66].

Griffi:1966:CBR

[GV66] D. S. Griffi and R. S. Varga. Comments on Bueckner’s “Remark on a Paper by D. S. Griffin and R. S. Varga” [SIAM J. Appl. Math. 14 (1966), no. 2, 414–414]. SIAM Journal on Applied Mathematics, 14(2):415–416, March 1966. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gautschi:1983:EBG

[GV83] Walter Gautschi and Richard S. Varga. Error bounds for Gaussian quadrature of analytic functions. SIAM Journal on Numerical Analysis, 20(6):1170–1186, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_141.pdf.

Golub:1988:OPH

[GV88] Gene H. Golub and Richard S. Varga. Obituary to Peter Henrici (13. September 1923–13. March 1987). Numerische Mathematik, 52(5):481–482, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Garabedian:1958:RRR

[GVB58] H. L. Garabedian, R. S. Varga, and G. G. Bilodeau. Reactor response to reactivity changes during a xenon transient. Nuclear Science and Engineering, 3(5):548–572, ???? 1958. CODEN NSE-NAO. ISSN 0029-5639 (print), 1943-748X (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_6.pdf.

Gerstenhaber:1953:APS

[GVC53] Murray Gerstenhaber, R. S. Varga, and James Clunie. Advanced problems and solutions: 4484. American Mathematical Monthly, 60(6):425, June/July 1953. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SGV^{+}52].

Henrici:1977:NTS

[Hen77] P. Henrici. Note on a theorem of Saff and Varga. In Padé and rational approximation (Proceedings of the International Symposium, Universit of South Florida, Tampa, Florida, 1976), pages 157–161. Academic Press, New York, NY, USA, 1977.

Hadjidimos:1988:NSC

[HN88] A. Hadjidimos and M. Neumann. A note on the SSOR convergence domain due to Neumaier and Varga. Linear Algebra and its Applications, 107(??):207–217, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902467.

Proceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Herbold:1969:EQE

[HSV69] R. J. Herbold, M. H. Schultz, and R. S. Varga. The effect of quadrature errors in the numerical solution of boundary value problems by variational techniques. Aequationes Mathematicae, 3(1–2):247–270, February 1969. CODEN AEMABN. ISSN 0001-9054 (print), 1420-

8903 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_54.pdf.

Holladay:1958:PNN

[HV58] John C. Holladay and Richard S. Varga. On powers of non-negative matrices. Proceedings of the American Mathematical Society, 9 (4):631–634, August 1958. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_7.pdf.

Hageman:1964:BIM

[HV64] Louis A. Hageman and Richard S. Varga. Block iterative methods for cyclically reduced matrix equations. Numerische Mathematik, 6 (1):106–119, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_25.pdf.

Hoffman:1970:PDG

[HV70] Alan J. Hoffman and Richard S. Varga. Patterns of dependence in generalizations of Gerschgorin’s theorem. SIAM Journal on Numerical Analysis, 7(4):571–574, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_62.pdf.

Hedstrom:1971:ABS

[HV71a] Gerald W. Hedstrom and Richard S. Varga. Application of Besov spaces to spline approximation. Journal of Approximation Theory, 4(3):295–327, September 1971. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_65.pdf.

Herbold:1971:EQE

[HV71b] R. J. Herbold and R. S. Varga. The effect of quadrature errors in the numerical solution of two-dimensional boundary value problems by variational techniques. Aequationes Mathematicae, 7(1):36–58, February 1971. CODEN AEMABN. ISSN 0001-9054 (print), 1420-

8903 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_68.pdf.

Householder:1970:NGI

[HVW70] Alston S. Householder, Richard S. Varga, and James H. Wilkinson. A note on Gerschgorin’s inclusion theorem for eigenvalues of matrices. Numerische Mathematik, 16(2):141–144, November 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://www.math.kent.edu/~varga/pub/paper_60.pdf.

Jakimovski:2004:MAS

[JSV^{+}04] Amnon Jakimovski, József Szabados, Richard S. Varga, Jean Sidon, T. N. T. Goodman, A. Meir, Seng Luan Lee, A. Cavaretta, Marcel G. de Bruin, Sherman D. Riemenschneider, Rong-Qing Jia, and N. Sivakumar. In memoriam: Ambikeshwar Sharma (1920–2003).

Journal of Approximation Theory, 131(1):1–18, 2004. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Jerome:1969:GSF

[JV69] J. W. Jerome and R. S. Varga. Generalizations of spline functions and applications to nonlinear boundary value and eigenvalue problems. In T. N. E. Greville, editor, Theory and Applications of Spline Functions (Proceedings of Seminar, Math. Research Center, Univ. of Wisconsin, Madison, Wis., 1968), pages 103–155. Academic Press, New York, NY, USA, 1969. URL http://www.math.

kent.edu/~varga/pub/paper_47.pdf.

Kostic:2009:GTL

[KCV09] V. Kostić, L. J. Cvetković, and R. S. Varga. Geršgorin-type localizations of generalized eigenvalues. Numerical Linear Algebra with Applications, 16(11–12):883–898, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_241.pdf.

Kincaid:1990:IML

[KH90] David R. Kincaid and Linda J. Hayes, editors. Iterative methods for large linear systems (Austin, TX, October 19–21, 1988). Academic Press, New York, NY, USA, 1990. ISBN 0-12-407475-8. LCCN QA432 .I84 1990. Papers from a conference held Oct. 19–21, 1988, at the Center for Numerical Analysis of the University of Texas at Austin. Banquet in honour of David M. Young’s 65th birthday.

Kolotilina:2005:BRR

[Kol05] Lily Kolotilina. Book review: Richard S. Varga, Geršgorin and his circles, Springer series in computational mathematics vol. 36, Springer-Verlag, Berlin, Heidelberg, New York (2004) ISBN-10 3-540-21100-4, x + 226 pp. Linear Algebra and its Applications, 402 (1):397–402, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kammerer:1974:QIS

[KRV74] W. J. Kammerer, G. W. Reddien, and R. S. Varga. Quadratic interpolatory splines. Numerische Mathematik, 22(4):241–259, August 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_85.pdf.

Kroo:1997:WPA

[KSV97] A. Kroó, J. Szabados, and R. S. Varga. Weighted polynomial approximation of some entire functions on the real line. Annals of Numerical Mathematics, 4(1–4):405–413, 1997. ISSN 1021-

2655. URL http://www.math.kent.edu/~varga/pub/paper_219.

pdf. The heritage of P. L. Chebyshev: a Festschrift in honor of the 70th birthday of T. J. Rivlin.

Kammerer:1972:ABN

[KV72] W. J. Kammerer and R. S. Varga. On asymptotically best norms for powers of operators. Numerische Mathematik, 20(2):93–98, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_74.pdf.

Krautstengl:1995:MGS

[KV95] Alan Krautstengl and Richard S. Varga. Minimal Gerschgorin sets for partitioned matrices. II. The spectral conjecture. Electronic Transactions on Numerical Analysis, 3:66–82, June 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.3.1995/pp66-82.dir/pp66-82.html;
 http://etna.mcs.kent.edu/vol.3.1995/pp66-82.dir/pp66-
82.pdf.

Kostic:2012:LGE

[KVC12] V. Kostić, R. S. Varga, and L. Cvetković. Localization of generalized eigenvalues by Cartesian ovals. Numerical Linear Algebra with Applications, 19(4):728–741, August 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kincaid:2010:LTD

[KVW10] David R. Kincaid, Richard S. Varga, and Charles H. Warlick. The life and times of Dr. David M. Young, Jr. Numerical Linear Algebra with Applications, 17(5):743–757, October 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ling:1975:API

[LRV75] William H. Ling, John A. Roulier, and Richard S. Varga. On approximation by polynomials increasing to the right of the interval. Journal of Approximation Theory, 14(4):285–295, 1975. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_87.pdf.

Loughry:1981:BCD

[LSH^{+}81] C. W. Loughry, D. B. Sheffer, R. H. Hamor, R. E. Herron, R. A. Liebelt, F. Proietti-Orlandi, and R. S. Varga. Breast cancer detection utilizing biostereometric analysis. Cancer Detection and Prevention, 4(??):589–594, ???? 1981. CODEN CDPRD4. ISSN 0361-090X (print), 1873-443X (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_136.pdf.

Lachance:1979:BIP

[LSV79a] M. Lachance, E. B. Saff, and R. S. Varga. Bounds for incomplete polynomials vanishing at both endpoints of an interval. In C. V. Coffman and G. J. Fix, editors, Constructive approaches to mathematical models: proceedings of a conference in honor of R. J. Duffin, Pittsburgh, PA, 10–14 July, 1978, pages 421–437. Academic Press, New York, NY, USA, 1979. ISBN 0-12-178150-X.

LCCN QA401 .C66. URL http://www.math.kent.edu/~varga/
pub/paper_117.pdf.

Lachance:1979:IPP

[LSV79b] Michael Lachance, Edward B. Saff, and Richard S. Varga. Inequalities for polynomials with a prescribed zero. Mathematische Zeitschrift, 168(2):105–116, 1979. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_113.pdf.

Levinger:1966:MGS

[LV66a] B. W. Levinger and R. S. Varga. Minimal Gerschgorin sets. II. Pacific Journal of Mathematics, 17(2):199–210, 1966. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL

http://projecteuclid.org/euclid.pjm/1102994613;   http:
//www.math.kent.edu/~varga/pub/paper_31.pdf.

Levinger:1966:POT

[LV66b] Bernard W. Levinger and Richard S. Varga. On a problem of O. Taussky. Pacific Journal of Mathematics, 19(3):473–487, July 1966. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

URL http://projecteuclid.org/euclid.pjm/1102993718;
http://www.math.kent.edu/~varga/pub/paper_34.pdf.

Li:1989:NSU

[LV89] Xiezhang Li and Richard S. Varga. A note on the SSOR and US-SOR iterative methods applied to p-cyclic matrices. Numerische Mathematik, 56(2–3):109–121, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_172.pdf.

Li:1990:NSU

[LV90] Xie Zhang Li and Richard S. Varga. A note on the SSOR and US-SOR iterative methods applied to p-cyclic matrices. In David R. Kincaid and Linda J. Hayes, editors, Iterative Methods for Large Linear Systems (Austin, TX, 1988), pages 235–249. Academic Press, New York, NY, USA, 1990. ISBN 0-12-407475-8. LCCN QA432 .I84 1990.

Marek:1967:IDA

[Mar67] Ivo Marek. An infinite dimensional analogue of R. S. Varga’s lemma. Commentationes Mathematicae Universitatis Carolinae, 8 (1):27–38, 1967. ISSN 0010-2628 (print), 1213-7243 (electronic).

Meinardus:1971:CTE

[MRTV71] G. Meinardus, A. R. Reddy, G. D. Taylor, and R. S. Varga. Converse theorems and extensions in Chebyshev rational approximation to certain entire functions in [0, +∞). Bulletin of the American Mathematical Society, 77:460–461, 1971. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_64.pdf.

Meinardus:1972:CTE

[MRTV72] G. Meinardus, A. R. Reddy, G. D. Taylor, and R. S. Varga. Converse theorems and extensions in Chebyshev rational approximation to certain entire functions in [∗ ∗ ∗w(∗ ∗ 0, +∞) ∗ ∗ ∗ w ∗ ∗. Transactions of the American Mathematical Society, 170(443):171–185, August 1972. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-

6850 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_71.pdf.

Moak:1979:ZJP

[MSV79] D. S. Moak, E. B. Saff, and R. S. Varga. On the zeros of Jacobi polynomials P_{n}^{(α_{n},β_{n})}(x). Transactions of the American Mathematical Society, 249(1):159–162, 1979. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_111.pdf.

Manteuffel:1995:ASI

[MSV95] Thomas A. Manteuffel, Gerhard Starke, and Richard S. Varga. Adaptive k-step iterative methods for nonsymmetric systems of linear equations. Electronic Transactions on Numerical Analysis, 3: 50–65, June 1995. CODEN ???? ISSN 1068-9613 (print), 1097-

4067 (electronic). URL http://etna.mcs.kent.edu/vol.3.1995/
pp50-65.dir/pp50-65.pdf.

Medley:1968:SIGb

[MV68a] Helen I. Medley and Richard S. Varga. On smallest isolated Gerschgorin disks for eigenvalues. II. Numerische Mathematik, 11 (4):320–323, May 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_38.pdf.

Medley:1968:SIGc

[MV68b] Helen I. Medley and Richard S. Varga. On smallest isolated Gerschgorin disks for eigenvalues. III. Numerische Mathematik, 11 (5):361–369, June 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_40.pdf.

Marek:1969:NBS

[MV69] Ivo Marek and Richard S. Varga. Nested bounds for the spectral radius. Numerische Mathematik, 14(1):49–70, December 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://www.math.kent.edu/~varga/pub/paper_53.pdf.

Meinardus:1970:CRA

[MV70] Günter Meinardus and Richard S. Varga. Chebyshev rational approximations to certain entire functions in [0, +∞). Journal of Approximation Theory, 3(3):300–309, September 1970. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_58.pdf.

Micchelli:2009:MTJ

[MV09] Charles A. Micchelli and Richard S. Varga. In memoriam: Theodore J. Rivlin (1926–2006). Journal of Approximation Theory, 161(1):1–8, 2009. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Niethammer:1984:CBI

[NdPV84] W. Niethammer, J. de Pillis, and R. S. Varga. Convergence of block iterative methods applied to sparse least-squares problems. Linear Algebra and its Applications, 58(??):327–341, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_143.pdf;   http://www.sciencedirect.com/science/
article/pii/0024379584902180.

Norfolk:1992:LBB

[NRV92] T. S. Norfolk, A. Ruttan, and R. S. Varga. A lower bound for the de Bruijn–Newman constant Λ. II. In A. A. Gonchar and E. B. Saff, editors, Progress in approximation theory (Tampa, FL, 1990), volume 19 of Springer Ser. Comput. Math., pages 403–418. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK /

etc., 1992. URL http://www.math.kent.edu/~varga/pub/paper_
193.pdf.

Neumann:1980:SSU

[NV80] Michael Neumann and Richard S. Varga. On the sharpness of some upper bounds for the spectral radii of S.O.R. iteration matrices. Numerische Mathematik, 35(1):69–79, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_119.pdf.

Niethammer:1983:ASI

[NV83] W. Niethammer and R. S. Varga. The analysis of k-step iterative methods for linear systems from summability theory. Numerische Mathematik, 41(2):177–206, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_138.pdf.

Neumaier:1984:ECD

[NV84] A. Neumaier and R. S. Varga. Exact convergence and divergence domains for the symmetric successive overrelaxation iterative (SSOR) method applied to H-matrices. Linear Algebra and its Applications, 58(??):261–272, April 1984. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_142.pdf;   http://
www.sciencedirect.com/science/article/pii/0024379584902167.

Niethammer:1989:RMN

[NV89] W. Niethammer and R. S. Varga. Relaxation methods for non-Hermitian linear systems. Results in Mathematics = Resultate der Mathematik, 16(3–4):308–320, 1989. ISSN 0378-6218. URL http:

//www.math.kent.edu/~varga/pub/paper_173.pdf.

Nabben:1994:LAP

[NV94] Reinhard Nabben and Richard S. Varga. A linear algebra proof that the inverse of a strictly ultrametric matrix is a strictly diagonally dominant Stieltjes matrix. SIAM Journal on Matrix Analysis and Applications, 15(1):107–113, January 1994. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.org/sam-bin/dbq/article/22823;
http://www.math.kent.edu/~varga/pub/paper_205.pdf.

Nabben:1995:GUM

[NV95a] Reinhard Nabben and Richard S. Varga. Generalized ultrametric matrices — a class of inverse M-matrices. Linear Algebra and its Applications, 220(1–3):365–390, April 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/laa/cas_sub/browse/browse.cgi?year=1995&
volume=220&issue=1-3&aid=9400086;   http://www.math.kent.
edu/~varga/pub/paper_212.pdf. Proceedings of the Workshop

“Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Nabben:1995:CIM

[NV95b] Reinhard Nabben and Richard S. Varga. On classes of inverse Zmatrices. Linear Algebra and its Applications, 223/224(1–3):521– 552, July ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print),

1873-1856 (electronic). URL http://www.elsevier.com/cgi-
bin/cas/tree/store/laa/cas_sub/browse/browse.cgi?year=
1995&volume=223-224&issue=1-3&aid=9592718;   http://www.
math.kent.edu/~varga/pub/paper_213.pdf. Special issue honor-

ing Miroslav Fiedler and Vlastimil Pták.

Noutsos:2012:PFT

[NV12] Dimitrios Noutsos and Richard S. Varga. On the Perron– Frobenius theory for complex matrices. Linear Algebra and its Applications, 437(4):1071–1088, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002437951200256X.

Olivier:1988:NPS

[ORV88] P. Olivier, Q. I. Rahman, and R. S. Varga. On a new proof and sharpenings of a result of Fejér on bounded partial sums. Linear Algebra and its Applications, 107(??):237–251, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_164.pdf;   http://www.sciencedirect.com/science/
article/pii/0024379588902480. Proceedings of the Victoria

Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Price:1968:NMH

[PCV68] H. S. Price, J. C. Cavendish, and R. S. Varga. Numerical methods of higher-order accuracy for diffusion–convection equations. Society of Petroleum Engineers journal, 8(??):293–303, September 1968. CODEN SPTJAJ. ISSN 0037-9999. URL http://www.math.kent.

edu/~varga/pub/paper_43.pdf.

Proietti-Orlandi:1988:BAB

[POVS^{+}88] F. Proietti-Orlandi, R. S. Varga, D. B. Sheffer, T. E. Price, and C. W. Loughry. Biostereometric analysis for breast cancer detection. Journal of Biomedical Engineering, 10(??):237–245, ???? 1988. CODEN JBIEDR. ISSN 0141-5425 (print), 2213-9478 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
163.pdf.

Perrin:1969:HON

[PPV69] F. M. Perrin, H. S. Price, and R. S. Varga. On higher-order numerical methods for nonlinear two-point boundary value problems. Numerische Mathematik, 13(2):180–198, May 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_49.pdf.

Price:1970:EBS

[PV70] Harvey S. Price and Richard S. Varga. Error bounds for semidiscrete Galerkin approximations of parabolic problems with applications to petroleum reservoir mechanics. In Numerical Solution of Field Problems in Continuum Physics (Proc. Sympos. Appl. Math., Durham, N.C., 1968), SIAM-AMS Proc., Vol. II, pages 74–94. American Mathematical Society, Providence, RI, USA, 1970. URL

http://www.math.kent.edu/~varga/pub/paper_56.pdf.

Pierce:1972:HOCa

[PV72a] J. G. Pierce and R. S. Varga. Higher order convergence results for the Rayleigh–Ritz method applied to eigenvalue problems. I: Estimates relating Rayleigh–Ritz and Galerkin approximations to eigenfunctions. SIAM Journal on Numerical Analysis, 9(1):137– 151, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-

7170 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_67.pdf.

Pierce:1972:HOCb

[PV72b] J. G. Pierce and R. S. Varga. Higher order convergence results for the Rayleigh–Ritz method applied to eigenvalue problems. II. Improved error bounds for eigenfunctions. Numerische Mathematik, 19(2):155–169, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_69.pdf.

Pritsker:1996:BSF

[PV96] Igor E. Pritsker and Richard S. Varga. Boundary singularities of Faber and Fourier series. Analysis, 16(3):283–295, 1996. ISSN 0174-

4747. URL http://www.math.kent.edu/~varga/pub/paper_217.
pdf.

Pritsker:1997:SCZ

[PV97a] Igor E. Pritsker and Richard S. Varga. The Szegő curve, zero distribution and weighted approximation. Transactions of the American Mathematical Society, 349(10):4085–4105, 1997. CODEN TAM-TAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_221.pdf.

Pritsker:1997:WPA

[PV97b] Igor E. Pritsker and Richard S. Varga. Weighted polynomial approximation in the complex plane. Electronic Research Announcements of the American Mathematical Society, 3:38–44, 1997.

ISSN 1079-6762. URL http://www.math.kent.edu/~varga/pub/
paper_220.pdf.

Pritsker:1998:WPA

[PV98a] I. E. Pritsker and R. S. Varga. Weighted polynomial approximation in the complex plane. Constructive Approximation, 14(4):475–492, 1998. ISSN 0176-4276 (print), 1432-0940 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_222.pdf.

Pritsker:1998:ZDS

[PV98b] I. E. Pritsker and R. S. Varga. Zero distribution, the Szegő curve, and weighted polynomial approximation in the complex plane. In J. W. Jerome, editor, Modelling and computation for applications in mathematics, science, and engineering (Evanston, IL, 1996), Numer. Math. Sci. Comput., pages 167–188. Oxford University Press,

Oxford, UK, 1998. URL http://www.math.kent.edu/~varga/
pub/paper_223.pdf.

Pritsker:1999:WRA

[PV99a] I. E. Pritsker and R. S. Varga. Weighted rational approximation in the complex plane. Journal de Mathématiques Pures et Appliquées, 78(2):177–202, 1999. CODEN JMPAAM. ISSN 0021-7824 (print), 1776-3371 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_224.pdf.

Pritsker:1999:RAV

[PV99b] Igor E. Pritsker and Richard S. Varga. Rational approximation with varying weights in the complex plane. In N. Papamichael, St. Ruscheweyh, and E. B. Saff, editors, Computational methods and function theory 1997 (Nicosia), volume 11 of Ser. Approx. Decompos., pages 437–448. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999. URL http://www.

math.kent.edu/~varga/pub/paper_226.pdf.

Price:1966:AOM

[PVW66] Harvey S. Price, Richard S. Varga, and Joseph E. Warren. Application of oscillation matrices to diffusion-convection equations. Journal of Mathematical Physics, 45:301–311, 1966. CODEN JMAPAQ.

ISSN 0097-1421. URL http://www.math.kent.edu/~varga/pub/
paper_33.pdf.

Reichel:1993:NLA

[RRV93] Lothar Reichel, Arden Ruttan, and Richard S. Varga, editors. Numerical linear algebra: proceedings of the Conference in Numerical Linear Algebra and Scientific Computation, Kent (Ohio), USA, March 13-=14, 1992. Walter de Gruyter, Berlin, Germany, 1993. ISBN 3-11-013784-4. LCCN QA184.C655 1993.

Ruscheweyh:1990:CMF

[RSSV90] Stephan Ruscheweyh, Edward B. Saff, Luis C. Salinas, and Richard S. Varga, editors. Computational Methods and Function Theory: Proceedings of the International Conference held at the Universidad Técnica Federico Santa Marí a, Valparaí so, March 13–18, 1989, volume 1435 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

LCCN QA3 .L28 no. 1435. URL http://link.springer.com/
book/10.1007/BFb0087892;   http://www.springerlink.com/
content/978-3-540-47139-4.

Rigler:1989:SLB

[RTV89] A. K. Rigler, S. Y. Trimble, and R. S. Varga. Sharp lower bounds for a generalized Jensen inequality. Rocky Mountain Journal of Mathematics, 19(1):353–373, 1989. CODEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic). URL http://www.math.kent.edu/ ~varga/pub/paper_170.pdf. Constructive Function Theory—86 Conference (Edmonton, AB, 1986).

Roulier:1978:APC

[RV78] John A. Roulier and Richard S. Varga. Another property of Chebyshev polynomials. Journal of Approximation Theory, 22(3):233– 242, 1978. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-

0430 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_105.pdf.

Ruscheweyh:1984:MMN

[RV84] S. Ruscheweyh and R. S. Varga. On the minimum moduli of normalized polynomials. Lecture Notes in Mathematics, 1105:150–159, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072408/;   http://www.math.kent.edu/~varga/
pub/paper_147.pdf.

Ruscheweyh:1986:MMN

[RV86] Stephan Ruscheweyh and Richard S. Varga. On the minimum moduli of normalized polynomials with two prescribed values.

Constructive Approximation, 2(4):349–368, 1986. ISSN 0176-4276 (print), 1432-0940 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_158.pdf.

Rodrigue:1988:CRE

[RV88] Garry Rodrigue and Richard Varga. Convergence rate estimates for iterative solutions of the biharmonic equation. Journal of Computational and Applied Mathematics, 24(1–2):129–146, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-

1778 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_167.pdf;   http://www.sciencedirect.com/science/
article/pii/0377042788903482. Iterative methods for the so-

lution of linear systems.

Ruttan:1989:RVC

[RV89a] Arden Ruttan and Richard S. Varga. Real vs. complex rational Chebyshev approximation on an interval. Rocky Mountain Journal of Mathematics, 19(1):375–381, 1989. CODEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_171.pdf. Constructive Func-

tion Theory—86 Conference (Edmonton, AB, 1986).

Ruttan:1989:UTR

[RV89b] Arden Ruttan and Richard S. Varga. A unified theory for real vs. complex rational Chebyshev approximation on an interval. Transactions of the American Mathematical Society, 312(2):681– 697, 1989. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-

6850 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_168.pdf.

Spiegel:1952:APSb

[SGV^{+}52] M. R. Spiegel, Murray Gerstenhaber, R. S. Varga, Simon T. Kao, D. J. Newman, and Paul Erdős. Advanced problems and solutions: Problems for solution: 4483–4487. American Mathematical Monthly, 59(4):254–255, April 1952. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sheffer:1986:VRB

[SPL^{+}86] D. B. Sheffer, T. E. Price, C. W. Loughry, B. L. Bolyard, W. M. Morek, and R. S. Varga. Validity and reliability of biostereometric measurement of the female breast. Annals of Biomedical Engineering, 14(1):1–14, January 1986. CODEN ABMECF. ISSN 0090-6964 (print), 1573-9686 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_155.pdf.

Saff:1976:EGC

[SSV76a] E. B. Saff, A. Schönhage, and R. S. Varga. Errata: “Geometric convergence to e^{−z} by rational functions with real poles” (Numer. Math. 25 (1975/76), no. 3, 307–322). Numerische Mathematik, 26 (2):226, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [SVN76].

Saff:1976:GCRb

[SSV76b] E. B. Saff, A. Schönhage, and R. S. Varga. Geometric convergence to e^{−z} by rational functions with real poles. Numerische Mathematik, 25(3):307–322, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_94.pdf. See errata [SSV76a].

Saff:1981:ERF

[SSV81] E. B. Saff, A. Sharma, and R. S. Varga. An extension to rational functions of a theorem of J. L. Walsh on differences of interpolating polynomials. RAIRO Analyse numérique/Numerical Analysis, 15 (4):371–390, 1981. ISSN 0399-0516. URL http://www.math.kent.

edu/~varga/pub/paper_133.pdf.

Sharma:1988:PLT

[SSV88] A. Sharma, J. Szabados, and R. S. Varga. 2-periodic lacunary trigonometric interpolation: the (0; M) case. In Blagovest Sendov et al., editors, Constructive theory of functions: proceedings of the International Conference on Constructive Theory of Functions, Varna, May 24–May 31, 1987, pages 420–427. Publishing House of the Bulgarian Academy of Sciences, Sofia, Bulgaria, 1988. LCCN

QA331 .I573 1987. URL http://www.math.kent.edu/~varga/
pub/paper_165.pdf.

Sharma:1991:SPT

[SSV91] A. Sharma, J. Szabados, and R. S. Varga. Some 2-periodic trigonometric interpolation problems on equidistant nodes. Analysis, 11(2–3):165–190, 1991. ISSN 0174-4747. URL http:

//www.math.kent.edu/~varga/pub/paper_186.pdf;   http://
www.math.kent.edu/~varga/pub/paper_187.pdf.

Sharma:1994:SPT

[SSV94] A. Sharma, J. Szabados, and R. S. Varga. Some 2-periodic trigonometric interpolation problems on equidistant nodes. II. Convergence. Studia Scientiarum Mathematicarum Hungarica, 29(3–4): 415–432, 1994. ISSN 0081-6906. URL http://www.math.kent.

edu/~varga/pub/paper_211.pdf.

Saff:1975:GCR

[SSV76] E. B. Saff, A. Schönhage, and R. S. Varga. Geometric convergence to e^{−z} by rational functions with real poles. Numerische Mathematik, 25(3):307–322, 1975/76. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saff:1980:IPE

[SUV80] E. B. Saff, J. L. Ullman, and R. S. Varga. Incomplete polynomials: an electrostatics approach. In E. W. (Elliott Ward) Cheney, editor, Approximation theory III: proceedings of a conference honoring Professor George G. Lorentz, held at the University of Texas at Austin, January 8–12, 1980, pages 769–782. Academic Press, New York, NY, USA, 1980. LCCN QA297.5 .A67. URL http://

www.math.kent.edu/~varga/pub/paper_122.pdf.

Schultz:1967:S

[SV67] M. H. Schultz and R. S. Varga. L-splines. Numerische Mathematik, 10(4):345–369, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_36.pdf.

Swartz:1972:EBS

[SV72] Blair K. Swartz and Richard S. Varga. Error bounds for spline and L-spline interpolation. Journal of Approximation Theory, 6(2):6– 49, July 1972. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-

0430 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_70.pdf.

Swartz:1973:NLI

[SV73] Blair K. Swartz and Richard S. Varga. A note on lacunary interpolation by splines. SIAM Journal on Numerical Analysis, 10(3):443– 447, June 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-

7170 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_79.pdf.

Saff:1975:CPA

[SV75a] E. B. Saff and R. S. Varga. Convergence of Padé approximants to e^{−z} on unbounded sets. Journal of Approximation Theory, 13(4):470–488, April 1975. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_86.pdf.

Saff:1975:ZPP

[SV75b] E. B. Saff and R. S. Varga. On the zeros and poles of Padé approximants to e^{z}. Numerische Mathematik, 25(1):1–14, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_

88.pdf. See erratum [SV76c].

Saff:1976:AOR

[SV76a] E. B. Saff and R. S. Varga. Angular overconvergence for rational functions converging geometrically on [0, +∞]. In A. G. Law and B. N. Sahney, editors, Theory of approximation, with applications (Proc. Conf., Univ. Calgary, Calgary, Alta., 1975; dedicated to the memory of Eckard Schmidt), pages 238–256. Academic Press, New

York, NY, USA, 1976. URL http://www.math.kent.edu/~varga/
pub/paper_89.pdf.

Saff:1976:BPT

[SV76b] E. B. Saff and R. S. Varga. The behavior of the Padé table for the exponential. In G. G. Lorentz, C. K. Chui, and L. L. Schumaker, editors, Approximation theory, II (Proc. Internat. Sympos., Univ. Texas, Austin, Tex., 1976), pages 519–531. ????, ????, 1976. URL

http://www.math.kent.edu/~varga/pub/paper_99.pdf.

Saff:1976:EZP

[SV76c] E. B. Saff and R. S. Varga. Erratum: “On the zeros and poles of Padé approximants to e^{z}” (Numer. Math. 25 (1975/76), no. 1, 1–14). Numerische Mathematik, 26(2):226, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [SV75b].

Saff:1976:GOR

[SV76d] E. B. Saff and R. S. Varga. Geometric overconvergence of rational functions in unbounded domains. Pacific Journal of Mathematics, 62(2):523–549, 1976. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http://projecteuclid.

org/euclid.pjm/1102867736;   http://www.math.kent.edu/~
varga/pub/paper_97.pdf.

Saff:1976:STC

[SV76e] E. B. Saff and R. S. Varga. On the sharpness of theorems concerning zero-free regions for certain sequences of polynomials. Numerische Mathematik, 26(4):345–354, December 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_98.pdf. See erratum [SV77a].

Saff:1976:ZFP

[SV76f] E. B. Saff and R. S. Varga. Zero-free parabolic regions for sequences of polynomials. SIAM Journal on Mathematical Analysis, 7(3):344– 357, May 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-

7154 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_91.pdf.

Saff:1977:EST

[SV77a] E. B. Saff and R. S. Varga. Erratum: “On the sharpness of theorems concerning zero-free regions for certain sequences of polynomials” (Numer. Math. 26 (1976), no. 4, 345–354). Numerische Mathematik, 28(1):119, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [SV76e].

Saff:1977:NBA

[SV77b] E. B. Saff and R. S. Varga. Nonuniqueness of best approximating complex rational functions. Bulletin of the American Mathematical Society, 83(3):375–377, 1977. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_101.pdf.

Saff:1977:ZPP

[SV77c] E. B. Saff and R. S. Varga. On the zeros and poles of Padé approximants to e^{z}. II. In E. B. Saff and R. S. Varga, editors, Padé and rational approximation: theory and applications: proceedings of an international symposium held at the University of South Florida, Tampa, Florida, December 15–17, 1976, pages 195–213. Academic Press, New York, NY, USA, 1977. URL http://www.math.kent.

edu/~varga/pub/paper_102.pdf.

Saff:1977:NLA

[SV77d] E. B. Saff and R. S. Varga, editors. Padé and Rational Approximation: Theory and Applications: Proceedings of an international symposium held at the University of South Florida, Tampa, Florida, December 15–17, 1976. Academic Press, New York, NY, USA, 1977. ISBN 0-12-614150-9. LCCN C20.7.P3 C66 1976.

Saff:1977:SOP

[SV77e] E. B. Saff and R. S. Varga. Some open problems concerning polynomials and rational functions. In E. B. Saff and R. S. Varga, editors, Padé and rational approximation (Proc. Internat. Sympos., Univ. South Florida, Tampa, Fla., 1976), pages 483–488. Academic Press,

New York, NY, USA, 1977. URL http://www.math.kent.edu/~
varga/pub/paper_103.pdf.

Saff:1978:NBC

[SV78a] E. B. Saff and R. S. Varga. Nonuniqueness of best complex rational approximations to real functions on real intervals. Journal of Approximation Theory, 23(1):78–85, 1978. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_106.pdf.

Saff:1978:IP

[SV78b] E. B. Saff and R. S. Varga. On incomplete polynomials. In ????, editor, Numerische Methoden der Approximationstheorie, Band 4 (Meeting, Math. Forschungsinst., Oberwolfach, 1977), volume 42 of Internat. Schriftenreihe Numer. Math., pages 281–298. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1978.

URL http://www.math.kent.edu/~varga/pub/paper_109.pdf.

Saff:1978:ZPP

[SV78c] E. B. Saff and R. S. Varga. On the zeros and poles of Padé approximants to e^{z}. III. Numerische Mathematik, 30(3):241–266, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_107.pdf.

Saff:1978:UAI

[SV78d] E. B. Saff and R. S. Varga. Uniform approximation by incomplete polynomials. International Journal of Mathematics and Mathematical Sciences, 1(4):407–420, 1978. CODEN ???? ISSN 0161-1712 (print), 1687-0425 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_110.pdf.

Saff:1979:SLT

[SV79] E. B. Saff and R. S. Varga. The sharpness of Lorentz’s theorem on incomplete polynomials. Transactions of the American Mathematical Society, 249(1):163–186, 1979. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_112.pdf. See corrigendum [].

Saff:1980:RSC

[SV80] E. B. Saff and R. S. Varga. Remarks on some conjectures of G. G. Lorentz. Journal of Approximation Theory, 30(1):29–36, September 1980. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-

0430 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_126.pdf.

Saff:1981:IPI

[SV81a] E. B. Saff and R. S. Varga. On incomplete polynomials. II. Pacific Journal of Mathematics, 92(1):161–172, 1981. CODEN PJ-MAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL

http://projecteuclid.org/euclid.pjm/1102737508;   http:
//www.math.kent.edu/~varga/pub/paper_129.pdf.

Saff:1981:LIP

[SV81b] Edward B. Saff and Richard S. Varga. On lacunary incomplete polynomials. Mathematische Zeitschrift, 177(3):297–314, 1981. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_130.pdf.

Szabados:1982:OCI

[SV82] J. Szabados and R. S. Varga. On the overconvergence of complex interpolating polynomials. Journal of Approximation Theory, 36(4): 346–363, 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-

0430 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_137.pdf.

Saff:1983:NSJ

[SV83a] E. B. Saff and R. S. Varga. A note on the sharpness of J. L. Walsh’s theorem and its extensions for interpolation in the roots of unity. Acta Mathematica Hungarica, 41(3–4):371–377, September 1983.

ISSN 0236-5294. URL http://www.math.kent.edu/~varga/pub/
paper_140.pdf.

Szabados:1983:OCI

[SV83b] J. Szabados and R. S. Varga. On the overconvergence of complex interpolating polynomials. II. Domain of geometric convergence to zero. Acta Scientiarum Mathematicarum (Szeged), 45(1-4):377–380, 1983. ISSN 0001-6969 (print), 2064-8316 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_139.pdf.

Sharma:1989:PLT

[SV89a] A. Sharma and R. S. Varga. On a 2-periodic lacunary trigonometric interpolation problem. In Charles K. Chui, Larry L. Schumaker, and Joseph D. Ward, editors, Approximation theory VI, Vol. II (College Station, TX, 1989), pages 585–588. Academic Press, New

York, NY, USA, 1989. URL http://www.math.kent.edu/~varga/
pub/paper_176.pdf.

Sharma:1989:PPL

[SV89b] A. Sharma and R. S. Varga. On a particular 2-periodic lacunary trigonometric interpolation problem on equidistant nodes. Results in Mathematics = Resultate der Mathematik, 16(3–4):383–404,

1989. ISSN 0378-6218. URL http://www.math.kent.edu/~varga/
pub/paper_174.pdf.

Starke:1993:HAF

[SV93] Gerhard Starke and Richard S. Varga. A hybrid Arnoldi–Faber iterative method for nonsymmetric systems of linear equations. Numerische Mathematik, 64(2):213–240, February 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_196.pdf.

Smithies:2006:SVD

[SV06] Laura Smithies and Richard S. Varga. Singular value decomposition Geršgorin sets. Linear Algebra and its Applications, 417(2–3): 370–380, September 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_236.pdf.

Saff:1976:GCRa

[SVN76] E. B. Saff, R. S. Varga, and W.-C. C. Ni. Geometric convergence of rational approximations to e^{−z} in infinite sectors. Numerische Mathematik, 26(2):211–225, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_92.pdf.

Taussky:1949:RTD

[Tau49] Olga Taussky. A recurring theorem on determinants. American Mathematical Monthly, 56(10):672–676, December 1949. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1972:PMH

[Tau72] Olga Taussky. Pairs of matrices having property L. In Richard S. Varga, editor, Gatlinburg V Symposium on Numerical Algebra, June 1972, page ?? ????, ????, 1972. URL http://www.math.

vt.edu/HHXX/history.html.

Todd:1999:HIU

[TV99] John Todd and Richard S. Varga. Hardy’s inequality and ultrametric matrices. Linear Algebra and its Applications, 302–303(1–3): 33–43, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.elsevier.nl/

gej-ng/10/30/19/117/25/28/abstract.html;   http://www.
elsevier.nl/gej-ng/10/30/19/117/25/28/article.pdf;   http:
//www.math.kent.edu/~varga/pub/paper_227.pdf. Special issue

dedicated to Hans Schneider (Madison, WI, 1998).

Varga:1952:SII

[Var52] Richard S. Varga. Semi-infinite and infinite strips free of zeros. Rendiconti del Seminario Matematico. Università e Politecnico Torino, 11:289–296, 1952. ISSN 0373-1243. URL http://www.math.kent.

edu/~varga/pub/paper_1.pdf.

Varga:1954:ECM

[Var54a] Richard S. Varga. Eigenvalues of circulant matrices. Pacific Journal of Mathematics, 4(1):151–160, 1954. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:/

/projecteuclid.org/euclid.pjm/1103044961;   http://www.
math.kent.edu/~varga/pub/paper_2.pdf.

Varga:1954:PSS

[Var54b] Richard S. Varga. Properties of a Special Set of Entire Functions and Their Respective Partial Sums. Ph.D. thesis, Harvard University, Cambridge, MA, USA, 1954.

Varga:1957:NST

[Var57a] R. S. Varga. Numerical solution of the two-group diffusion equation in x − y geometry. IRE Transactions on Nuclear Science, NS-4:

52–62, 1957. CODEN IRNSAM. ISSN 0096-2015 (print), 2374-

9814 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_4.pdf. December.

Varga:1957:CSO

[Var57b] Richard S. Varga. A comparison of the successive overrelaxation method and semi-iterative methods using Chebyshev polynomials. Journal of the Society for Industrial and Applied Mathematics, 5(2):39–46, June 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic). URL http://www.math.kent.

edu/~varga/pub/paper_5.pdf.

Varga:1959:OSO

[Var59a] Richard S. Varga. Orderings of the successive overrelaxation scheme. Pacific Journal of Mathematics, 9(3):925–939, Fall 1959. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

URL http://projecteuclid.org/euclid.pjm/1103039130;
http://www.math.kent.edu/~varga/pub/paper_13.pdf.

Varga:1959:CMG

[Var59b] Richard S. Varga. p-cyclic matrices: a generalization of the Young–Frankel successive overrelaxation scheme. Pacific Journal of Mathematics, 9(2):617–628, Summer 1959. CODEN PJ-MAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL

http://projecteuclid.org/euclid.pjm/1103039283;   http:
//www.math.kent.edu/~varga/pub/paper_11.pdf.

Varga:1960:RAI

[Var60a] R. S. Varga. Over-relaxation applied to implicit alternating direction methods. In ????, editor, Information Processing: Proceedings of the International Conference on Information Processing, UN-ESCO, Paris, France [15–20 June 1959], pages 85–90. UNESCO and R. Oldenbourg and Butterworths, Paris, France and Munich, Germany and London, UK, 1960. ISSN 0074-3127. LCCN QA76 .I578; Q300 .F60.

Varga:1960:FNI

[Var60b] Richard S. Varga. Factorization and normalized iterative methods. In ????, editor, Boundary Problems in Differential Equations: proceedings of a symposium conducted by the Mathematics Research Center at the University of Wisconsin, Madison, April 20–22, 1959, pages 121–142. University of Wisconsin Press, Madison, WI, USA,

1960. LCCN QA3 .U45 no. 2. URL http://www.math.kent.edu/
~varga/pub/paper_14.pdf.

Varga:1960:OAI

[Var60c] Richard S. Varga. Overrelaxation applied to implicit alternating direction methods. Information Processing, pages 342–??, 1960.

ISSN 0074-3127. URL http://www.math.kent.edu/~varga/pub/
paper_15.pdf.

Varga:1961:NMS

[Var61a] Richard S. Varga. Numerical methods for solving multi-dimensional multi-group diffusion equations. In Nuclear reactor theory. [Proceedings of the Eleventh Symposium in Applied Mathematics of the American Mathematics Society., volume XI, pages 164–189. American Mathematical Society, Providence, RI, USA, 1961. LCCN

C787.N8. URL http://www.math.kent.edu/~varga/pub/paper_
16.pdf.

Varga:1961:HOS

[Var61b] Richard S. Varga. On higher order stable implicit methods for solving parabolic partial differential equations. Journal of Mathematical Physics, 40:220–231, 1961. CODEN JMAPAQ. ISSN 0097-1421.

URL http://www.math.kent.edu/~varga/pub/paper_19.pdf.

Varga:1962:MIA

[Var62] Richard S. Varga. Matrix Iterative Analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1962. xiii + 322 pp. LCCN QA263 .V3 1962.

Varga:1963:VSO

[Var63] Richard S. Varga. On variants of successive overrelaxation and alternating direction implicit methods. Information Processing, ?? (??):203–204, ???? 1963. ISSN 0074-3127. URL http://www.math.

kent.edu/~varga/pub/paper_22.pdf.

Varga:1964:RPP

[Var64a] R. S. Varga. Recent publications and presentations: Numerical Analysis, by Zdenek Kopal. American Mathematical Monthly, 71(1):107, January 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Varga:1964:SIG

[Var64b] Richard S. Varga. On smallest isolated Gerschgorin disks for eigenvalues. Numerische Mathematik, 6(1):366–376, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
26.pdf.

Varga:1965:IMS

[Var65a] R. S. Varga. Iterative methods for solving matrix equations. American Mathematical Monthly, 72(2 (Part 2)):67–74, February 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_

27.pdf. Part 2: Computers and Computing.

Varga:1965:RPP

[Var65b] R. S. Varga. Recent publications and presentations: Numerical Analysis, by Nathaniel Macon. American Mathematical Monthly, 72(4):440, April 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Varga:1965:MGS

[Var65c] Richard S. Varga. Minimal Gerschgorin sets. Pacific Journal of Mathematics, 15(2):719–729, 1965. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:/

/projecteuclid.org/euclid.pjm/1102995820;   http://www.
math.kent.edu/~varga/pub/paper_29.pdf.

Varga:1966:BRT

[Var66a] R. S. Varga. Book review: The Theory of Matrices in Numerical Analysis (Alston S. Householder). SIAM Review, 8(1):127–128, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varga:1966:HIT

[Var66b] Richard S. Varga. Hermite interpolation-type Ritz methods for two-point boundary value problems. In J. H. (James H.) Bramble and Bert Hubbard, editors, Numerical Solution of Partial Differential Equations (Proceedings of a Symposium at the University of Maryland, 1965), pages 365–373. Academic Press, New York, NY,

USA, 1966. ISSN 0362-3017. URL http://www.math.kent.edu/~
varga/pub/paper_30.pdf.

Varga:1966:DMP

[Var66c] Richard S. Varga. On a discrete maximum principle. SIAM Journal on Numerical Analysis, 3(2):355–359, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_32.pdf.

Varga:1967:ASA

[Var67a] Richard S. Varga. Abstract: Some applications of numerical analysis to physical problems. SIAM Journal on Numerical Analysis, 4 (1):129, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varga:1967:BRI

[Var67b] Richard S. Varga. Book review: Iterative Solution of Elliptic Systems and Applications to the Neutron Diffusion Equations of Reactor Physics (Eugene L. Wachspress). SIAM Review, 9(4):756–758, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varga:1968:NPP

[Var68a] Richard S. Varga. Nonnegatively posed problems and completely monotonic functions. Linear Algebra and its Applications, 1(3): 329–347, July 1968. CODEN LAAPAW. ISSN 0024-3795 (print),

1873-1856 (electronic).  URL http://www.math.kent.edu/~
varga/pub/paper_44.pdf;   http://www.sciencedirect.com/
science/article/pii/002437956890013X.

Varga:1968:ERB

[Var68b] Richard S. Varga. On an extension of a result of S. N. Bernstein. Journal of Approximation Theory, 1(2):176–179, September 1968. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

URL http://www.math.kent.edu/~varga/pub/paper_46.pdf.

Varga:1969:EBS

[Var69] Richard S. Varga. Error bounds for spline interpolation. In I. J. Schoenberg, editor, Approximations, with special emphasis on spline functions; proceedings of a symposium conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, May 5–7, 1969, volume 23, pages 367–388. Academic Press, New York, NY, USA, 1969. ISBN 0-12-628850-X. LCCN QA3 .U45 no. 23. URL http://www.math.kent.

edu/~varga/pub/paper_52.pdf.

Varga:1970:ANMa

[Var70a] Richard S. Varga. Accurate numerical methods for nonlinear boundary value problems. In G. Birkhoff and R. S. Varga, editors, Studies in Numerical Analysis, 2: Numerical Solutions of Nonlinear Problems (Symposia, SIAM, Philadelphia, Pa., 1968), pages 99–113. Society for Industrial and Applied Mathematics, Philadel-

phia, PA, USA, 1970. URL http://www.math.kent.edu/~varga/
pub/paper_55.pdf.

Varga:1970:ANMb

[Var70b] Richard S. Varga. Accurate numerical methods for nonlinear boundary value problems. In Garrett Birkhoff and Richard S. Varga, editors, Numerical solution of field problems in continuum physics; proceedings of a symposium in applied mathematics [Durham, N.C., 1968], pages 152–167. American Mathematical Society, Providence, RI, USA, 1970. LCCN QC174.45.

Varga:1970:MGS

[Var70c] Richard S. Varga. Minimal Gerschgorin sets for partitioned matrices. SIAM Journal on Numerical Analysis, 7(4):493–507, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-

7170 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_61.pdf.

Varga:1971:FAA

[Var71a] R. S. Varga. Functional Analysis and Approximation Theory in Numerical Analysis. Conference Board of the Mathematical Sciences Regional Conference Series in Applied Mathematics, No. 3. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1971. v + 76 pp.

Varga:1971:SRA

[Var71b] Richard S. Varga. Some results in approximation theory with applications to numerical analysis. In B. E. Hubbard, editor, Numerical Solution of Partial differential Equations II. Academic Press, New

York, NY, USA, 1971. URL http://www.math.kent.edu/~varga/
pub/paper_63.pdf.

Varga:1972:CSD

[Var72a] Richard S. Varga. Chebyshev semi-discrete approximations for linear parabolic problems. In P. L. Butzer, J.-P. Kahane, and B. Sz.- Nagy, editors, Linear operators and approximation (Proc. Conf., Oberwolfach, 1971), volume 20 of Internat. Ser. Numer. Math., pages 452–460. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1972. URL http://www.math.kent.

edu/~varga/pub/paper_75.pdf.

Varga:1972:RIA

[Var72b] Richard S. Varga. The role of interpolation and approximation theory in variational and projectional methods for solving partial differential equations. In ????, editor, Information processing 71 (Proc. IFIP Congr., Ljubljana, 1971), Vol. 2: Applications, pages 1185–1190. North-Holland Publishing Co., Amsterdam, The Netherlands, 1972. URL http://www.math.kent.edu/

~varga/pub/paper_73.pdf.

Varga:1973:CSD

[Var73a] Richard S. Varga. Chebyshev semi-discrete approximations for linear parabolic problems. In A. Meir and A. Sharma, editors, Spline functions and approximation theory. Proceedings of the Symposium held at the University of Alberta, Edmonton, May 29 to June 1, 1972, volume 21 of International series of numerical mathematics, pages 383–386. ????, ????, 1973. ISBN 3-7643-0670-X. LCCN QA224 .S65.

Varga:1973:ESO

[Var73b] Richard S. Varga. Extensions of the successive overrelaxation theory with applications to finite element approximations. In J. J. H. Miller, editor, Topics in numerical analysis (Proc. Roy. Irish Acad. Conf., University College, Dublin, 1972), pages 329–343. Academic Press, New York, NY, USA, 1973. URL http://www.math.kent.

edu/~varga/pub/paper_81.pdf.

Varga:1973:CBI

[Var73c] Richard S. Varga. On a connection between infima of norms and eigenvalues of associated operators. Linear Algebra and its Applications, 6(??):249–256, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.math.kent.edu/~varga/pub/paper_77.pdf;   http://
www.sciencedirect.com/science/article/pii/0024379573900256.

Varga:1976:MTR

[Var76a] Richard S. Varga. M-matrix theory and recent results in numerical linear algebra. In James R. Bunch and Donald J. Rose, editors, Sparse matrix computations: proceedings of the Symposium on Sparse Matrix Computations at Argonne National Laboratory on September 9–11, 1975, pages 375–387. Academic Press, New York, NY, USA, 1976. ISBN 0-12-141050-1. LCCN QA188. URL

http://www.math.kent.edu/~varga/pub/paper_96.pdf.

Varga:1976:DDA

[Var76b] Richard S. Varga. On diagonal dominance arguments for bounding ||A^{−1}||_{∞}. Linear Algebra and its Applications, 14(3):211–217, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
95.pdf;   http://www.sciencedirect.com/science/article/
pii/0024379576900677.

Varga:1976:RTD

[Var76c] Richard S. Varga. On recurring theorems on diagonal dominance. Linear Algebra and its Applications, 13(1–2):1–9, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_
90.pdf;   http://www.sciencedirect.com/science/article/

pii/0024379576900379. Collection of articles dedicated to Olga Taussky Todd.

Varga:1977:NOQ

[Var77] Richard S. Varga. A note on an open question on ω- and τ-matrices. Linear Algebra and its Applications, 18(1):45–52, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_100.pdf;   http://www.sciencedirect.com/science/
article/pii/0024379577900787.

Varga:1978:RRL

[Var78] Ričard Varga. Recent results in linear algebra and their applications. In ????, editor, Numerical methods in linear algebra (Proc. Third Sem. Methods of Numer. Appl. Math., Novosibirsk, 1978) (Russian), pages 5–15. Akad. Nauk SSSR Sibirsk. Otdel., Vychisl. Tsentr, Novosibirsk, USSR, 1978. URL http://www.math.kent.

edu/~varga/pub/paper_114.pdf.

Varga:1982:TPR

[Var82] Richard S. Varga. Topics in Polynomial and Rational Interpolation and Approximation, volume 81 of Séminaire de Mathématiques Supérieures [Seminar on Higher Mathematics]. Presses de l’Université de Montréal, Montréal, Québec, Canada, 1982. ISBN 2-7606-0573-6. 136 pp.

Varga:1984:SRR

[Var84] Richard S. Varga. A survey of recent results on iterative methods for solving large sparse linear systems. In G. Birkhoff and A. Schoenstadt, editors, Elliptic problem solvers, II (Monterey, Calif., 1983), pages 197–217. Academic Press, New York, NY, USA, 1984. URL

http://www.math.kent.edu/~varga/pub/paper_145.pdf.

Varga:1986:SCS

[Var86] Richard S. Varga. Scientific computation on some mathematical conjectures. In Charles K. (Kam tai) Chui, Larry L. Shumaker, and Joseph D. Ward, editors, Approximation theory, V (College Station, Texas, 1986), pages 191–209. Academic Press, New York, NY, USA, 1986. ISBN 0-12-174581-3 (hardcover). LCCN QA221

.I56 1986. URL http://www.math.kent.edu/~varga/pub/paper_
159.pdf.

Varga:1988:SCS

[Var88] Richard S. Varga. Scientific computation on some mathematical conjectures. In ????, editor, Visiting scholars’ lectures—1987 (Lubbock, TX), volume 15 of Texas Tech Univ. Math. Ser., pages 119– 136. Texas Tech University, Lubbock, TX, USA, 1988.

Varga:1990:RUM

[Var90a] Richard S. Varga. Reminiscences on the University of Michigan summer schools, the Gatlinburg Symposia, and Numerische Mathematik. In Nash [Nas90], pages 206–210. ISBN 0-201-50814-1. LCCN

QA76.17 .H59 1990. URL http://www.math.kent.edu/~varga/
pub/paper_179.pdf.

Varga:1990:SCM

[Var90b] Richard S. Varga. Scientific Computation on Mathematical Problems and Conjectures, volume 60 of CBMS-NSF Regional Conference Series in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-257-2. vi + 122 pp.

Varga:1992:HHP

[Var92a] Richard S. Varga. How high-precision calculations can stimulate mathematical research. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4):177–193, September 1992.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.math.kent.edu/~varga/pub/paper_

190.pdf. A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Varga:1992:GMI

[Var92b] Richard S. Varga. On a generalization of Mahler’s inequality. Analysis, 12(3–4):319–333, 1992. ISSN 0174-4747. URL http://www.

math.kent.edu/~varga/pub/paper_191.pdf.

Varga:1992:SCMa

[Var92c] Richard S. Varga. Scientific computation on mathematical problems and conjectures [Chapters 1–3]. Alkalmazott Matematikai Lapok, 16(1–2):143–213, 1992. ISSN 0133-3399. Translated from the English original [MR1068317 (92b:65012)] by Jószef Szabados.

Varga:1992:SCMb

[Var92d] Richard S. Varga. Scientific computation on mathematical problems and conjectures [Chapters 4–6]. Alkalmazott Matematikai Lapok, 16(3–4):349–408, 1992. ISSN 0133-3399. Translated from the English original [MR1068317 (92b:65012)] by József Szabados.

Varga:1998:MGB

[Var98] Richard S. Varga. In memoriam: Garrett Birkhoff: January 10, 1911–November 11, 1996. Journal of Approximation Theory, 95(1): iv, 1–4, 1998. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Varga:2000:MIA

[Var00a] Richard S. Varga. Matrix Iterative Analysis, volume 27 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 3-642-05154-5 (paperback), 3-540-66321-5 (hardcover), 3-642-05156-1 (ebook). ISSN 0179-3632. x + 358 pp. LCCN QA297.8

.V37 2000. URL http://lcweb.loc.gov/catdir/toc/99050196.
html;   http://www.loc.gov/catdir/enhancements/fy0815/
99050196-d.html;   http://www.springerlink.com/openurl.
asp?genre=book&isbn=978-3-642-05161-6.

Varga:2000:R

[Var00b] Richard S. Varga. Reminiscences. Numerical Algorithms, 25(1-4): xv–xvi, 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Varga:2000:SCR

[Var00c] Richard S. Varga. Scientific computations related to the Riemann hypothesis. Sūrikaisekikenkyūsho Kōkyūroku, (1145):19–30, 2000. Numerical solution of partial differential equations and related topics (Japanese) (Kyoto, 1999).

Varga:2001:GDB

[Var01a] Richard S. Varga. Gerschgorin disks, Brauer ovals of Cassini (a vindication), and Brualdi sets. Information, 4(2):171–178, 2001.

ISSN 1343-4500. URL http://www.math.kent.edu/~varga/pub/
paper_232.pdf.

Varga:2001:GT

[Var01b] Richard S. Varga. Gerschgorin’s theorem. In ????, editor, Encyclopaedia of Mathematics, Supplement III, pages 180–181. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0198-3. URL http://www.math.

kent.edu/~varga/pub/paper_234.pdf.

Varga:2001:GTE

[Var01c] Richard S. Varga. Geršgorin-type eigenvalue inclusion theorems and their sharpness. Electronic Transactions on Numerical Analysis, 12:113–133, 2001. CODEN ???? ISSN 1068-9613 (print),

1097-4067 (electronic). URL http://etna.mcs.kent.edu/vol.
12.2001/pp113-133.dir/pp113-133.pdf.

Varga:2004:GHC

[Var04] Richard S. Varga. Geršgorin and His Circles, volume 36 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21100-4 (hardcover), 3-642-17798-0 (e-book). ISSN 0179-3632. x + 226 pp. LCCN QA193 .V37 2004. URL http://www.

loc.gov/catdir/enhancements/fy0815/2004104814-d.html.

Varga:2009:MIA

[Var09] Richard S. Varga. Matrix Iterative Analysis, volume 27 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second revised and expanded edition, 2009. ISBN 3-642-05154-5 (softcover), 3-642-05156-1 (e-book). ISSN 0179-3632 (print), 2198-3712 (electronic). x + 358 pp. LCCN QA263 .V3 2009.

Varga:1981:FM

[VC81] Richard S. Varga and Da Yong Cai. On the LU factorization of M-matrices. Numerische Mathematik, 38(2):179–192, December 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_132.pdf.

Varga:1982:FMC

[VC82] R. S. Varga and D. Y. Cai. On the LU factorization of M-matrices: cardinality of the set P_{n}^{g}(A). SIAM Journal on Algebraic and Discrete Methods, 3(2):250–259, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_135.pdf.

Varga:1985:BCA

[VC85] Richard S. Varga and Amos J. Carpenter. On the Bernstein conjecture in approximation theory. Constructive Approximation, 1 (4):333–348, 1985. ISSN 0176-4276 (print), 1432-0940 (electronic).

URL http://www.math.kent.edu/~varga/pub/paper_150.pdf.

Varga:1986:OOG

[VC86] R. S. Varga and A. J. Carpenter. Ob odnoi gipoteze S. Bernsteina v teorii priblizhenii. (Russian) [On one conjecture of S. Bernstein in approximation theory]. Matematicheskii sbornik, 129(4):535–548, ???? 1986. ISSN 1064-5616 (print), 1468-4802 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_154.pdf.

Varga:1987:OHB

[VC87] R. S. Varga and A. J. Carpenter. On a conjecture of S. Bernstein in approximation theory. Mathematics of the USSR. Sbornik, 57(2):547–560, ???? 1987. CODEN MUSBBS. ISSN 0025-

5734. URL http://www.math.kent.edu/~varga/pub/paper_154.

pdf. Reprint in English of original [VC85].

Varga:1990:AZP

[VC90] R. S. Varga and A. J. Carpenter. Asymptotics for the zeros of the partial sums of e^{z}. II. Lecture Notes in Mathematics, 1435:201–207, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087909/;   http://www.math.kent.edu/~varga/

pub/paper_178.pdf. Computational methods and function theory (Valparaíso, Chile, March 1989).

Varga:1992:SNR

[VC92] Richard S. Varga and Amos J. Carpenter. Some numerical results on best uniform rational approximation of x^{α} on [0, 1]. Numerical Algorithms, 2(2):171–185, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_189.pdf.

Varga:1994:AZP

[VC94] Richard S. Varga and Amos J. Carpenter. Asymptotics for the zeros and poles of normalized Padé approximants to e^{z}. Numerische Mathematik, 68(1):169–185, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.de/link/service/journals/00211/bibs/4068001/40680169.
htm;   http://science.springer.de/nmee/bibs/4068001/40680169.
htm;   http://www.math.kent.edu/~varga/pub/paper_207.pdf.

Varga:2000:ZPS

[VC00] Richard S. Varga and Amos J. Carpenter. Zeros of the partial sums of cos(z) and sin(z). I. Numerical Algorithms, 25(1– 4):363–375, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/3/2/abstract.htm;
 http://ipsapp007.kluweronline.com/content/getfile/5058/
3/2/fulltext.pdf;   http://www.math.kent.edu/~varga/pub/

paper_230.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Varga:2001:ZPS

[VC01] Richard S. Varga and Amos J. Carpenter. Zeros of the partial sums of cos(z) and sin(z). II. Numerische Mathematik, 90(2):371–400, December 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090002/10900371.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090002/10900371.pdf;   http://www.math.kent.edu/
~varga/pub/paper_233.pdf.

Varga:2010:ZPS

[VC10] Richard S. Varga and Amos J. Carpenter. Zeros of the partial sums of cos(z) and sin(z), III. Applied Numerical Mathematics: Transactions of IMACS, 60(4):298–313, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http://www.

math.kent.edu/~varga/pub/paper_242.pdf.

Varga:2008:AMG

[VCK08] Richard S. Varga, Ljiljana Cvetković, and Vladimir Kostić. Approximation of the minimal Geršgorin set of a square complex matrix. Electronic Transactions on Numerical Analysis, 30:398–405, 2008. CODEN ???? ISSN 1068-9613 (print),

1097-4067 (electronic). URL http://etna.mcs.kent.edu/vol.
30.2008/pp398-405.dir/pp398-405.pdf;   http://www.math.
kent.edu/~varga/pub/paper_240.pdf.

Varga:2008:DMZ

[VCL08] Richard S. Varga, Amos J. Carpenter, and Bryan W. Lewis. The dynamical motion of the zeros of the partial sums of e^{z}, and its relationship to discrepancy theory. Electronic Transactions on Numerical Analysis, 30:128–143, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.

mcs.kent.edu/vol.30.2008/pp128-143.dir/pp128-143.pdf;
http://www.math.kent.edu/~varga/pub/paper_239.pdf.

Varga:1986:CBA

[VK86] R. S. Varga and A. Dzh. Karpenter. A conjecture of S. Bernstein in approximation theory. Matematicheskiĭ Sbornik. Novaya Seriya, 129(171)(4):535–548, 591–592, 1986. ISSN 0368-8666.

Varga:1995:MGS

[VK95] Richard S. Varga and Alan Krautstengl. Minimal Gerschgorin sets for partitioned matrices. III. Sharpness of boundaries and monotonicity as a function of the partition. Electronic Transactions on Numerical Analysis, 3:83–95, September 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.mcs.

kent.edu/vol.3.1995/pp83-95.dir/pp83-95.pdf.

Varga:1999:GTP

[VK99] Richard S. Varga and Alan Krautstengl. On Geršgorin-type problems and ovals of Cassini. Electronic Transactions on Numerical Analysis, 8:15–20, 1999. CODEN ???? ISSN 1068-9613 (print),

1097-4067 (electronic). URL http://etna.mcs.kent.edu/vol.8.
1999/pp15-20.dir/pp15-20.pdf.

Varga:1973:MGS

[VL73] Richard S. Varga and Bernard W. Levinger. On minimal Gerschgorin sets for families of norms. Numerische Mathematik, 20 (3):252–256, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.math.kent.edu/

~varga/pub/paper_80.pdf.

Varga:1959:TNS

[VM59] R. S. Varga and M. A. Martino. The theory for the numerical solution of time-dependent and time-independent multigroup diffusion equations. In ????, editor, Proceedings of the Second United Nations International Conference on the Peaceful Uses of Atomic Energy (Geneva, Switzerland, 1–13 September 1958), volume 15, pages 570–577. Pergamon Press, London, UK, 1959. URL http:/

/www.math.kent.edu/~varga/pub/paper_12.pdf.

Varga:1993:ADI

[VN93a] Richard S. Varga and Reinhard Nabben. An algorithm for determining if the inverse of a strictly diagonally dominant Stieltjes matrix is strictly ultrametric. Numerische Mathematik, 65(4):493–501, August 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_200.pdf.

Varga:1993:SUM

[VN93b] Richard S. Varga and Reinhard Nabben. On symmetric ultrametric matrices. In L. Reichel, A. Ruttan, and R. S. Varga, editors, Numerical linear algebra (Kent, OH, 1992), pages 193–199. Walter de Gruyter, Berlin, Germany, 1993. URL http://www.math.kent.

edu/~varga/pub/paper_199.pdf.

Varga:1993:NPA

[VN93c] Richard S. Varga and Wilhelm Niethammer. A note on a perturbation analysis of iterative methods, with an application to the SSOR and ADI iterative method. In ????, editor, Proceedings of the seminar Numerical Mathematics in Theory and Practice, University of West Bohemia, 1993, pages 18–27. ????, ????, 1993. URL http:/

/www.math.kent.edu/~varga/pub/paper_203.pdf.

Varga:1984:CMS

[VNC84] R. S. Varga, W. Niethammer, and D.-Y. Cai. p-cyclic matrices and the symmetric successive overrelaxation method. Linear Algebra and its Applications, 58(??):425–439, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_144.pdf;   http://www.sciencedirect.com/science/
article/pii/0024379584902234.

Varga:1996:CTP

[VP96] Richard S. Varga and Igor E. Pritsker. On a counterexample in the theory of polynomials having concentration at low degrees. Analysis, 16(4):365–378, 1996. ISSN 0174-4747. URL http://www.math.

kent.edu/~varga/pub/paper_218.pdf.

Varga:1991:RVC

[VR91a] Richard S. Varga and Arden Ruttan. Real vs. complex best rational approximation. In Charles K. Chui, editor, Approximation theory and functional analysis (College Station, TX, 1990), pages 215– 236. Academic Press, New York, NY, USA, 1991. URL http://

www.math.kent.edu/~varga/pub/paper_182.pdf.

Verschelde:1991:BRN

[VR91b] J. Verschelde and D. Roose. Book review: Numerical Continuation Methods: an Introduction: E. L. Allgower, K. Georg, R. L. Graham, J. Stoer, R. Varga (Eds.) Springer-Verlag, 1990, Springer Series in Computational Mathematics. Volume 13. Approx. 400 pages, 37 figures. Hardcover, DM 128,-ISBN 3-540-12760-7. Journal of Computational and Applied Mathematics, 34(2):N2– N3, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037704279190050T.

Varga:2012:ANM

[VR12] Richard S. Varga and Alessandro Rizzo. An application of nonnegative matrices to the synchronization of chaotic oscillators. Linear Algebra and its Applications, 436(2):265–275, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.math.kent.edu/~varga/pub/
paper_243.pdf;   http://www.sciencedirect.com/science/
article/pii/S002437951000354X.

Varga:1991:NRB

[VRK91] R. S. Varga, A. Ruttan, and A. Dzh. Karpenter. Chislennie rezultati o nailuchshikh ravnomernikh ratsionalnikh approksimatsiyakh funktsii |x| na otrezke [−1, +1]. (Russian) [numerical results on the best uniform rational approximations of the function |x| on the interval [−1, 1]. Matematicheskii sbornik, 182(11):1523–1541, 1991. ISSN 0368-8666.

Varga:1993:NRB

[VRK93] R. S. Varga, A. Ruttan, and A. Dzh. Karpenter. Numerical results on the best uniform rational approximations of the function |x| on the interval [−1, 1]. Mathematics of the USSR. Sbornik, 74(2):271–290, ???? 1993. CODEN MUSBBS. ISSN 0025-5734.

URL http://www.math.kent.edu/~varga/pub/paper_197.pdf;
 https://www.turpion.org/php/paper.phtml?journal_id=sm&
paper_id=3347.

Varga:1980:IFM

[VSM80] R. S. Varga, E. B. Saff, and V. Mehrmann. Incomplete factorizations of matrices and connections with H-matrices. SIAM Journal on Numerical Analysis, 17(6):787–793, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL

http://www.math.kent.edu/~varga/pub/paper_123.pdf.

Varga:1974:CFN

[VV74] R. Varga and R. Varga. Funktsional′nyĭ analiz i teoriya approksimatsii v chislennom analize. Izdat. “Mir”, Moscow, USSR, 1974. 126 pp. Translated from the English by Ju. A. Kuznecov and A. M. Macokin, Edited by G. I. Marčuk.

Varga:1985:ROG

[VW85] Richard S. Varga and Wen Da Wu. On the rate of overconvergence of the generalized Eneström–Kakeya functional for polynomials. Journal of Computational Mathematics, 3(3):275–288, 1985. CO-DEN JCMMEB. ISSN 0254-9409 (print), 1991-7139 (electronic).

URL http://www.math.kent.edu/~varga/pub/paper_151.pdf.